### v2

# Association and Independence Test for Random Objects

Abstract
We develop a unified framework for testing independence and quantifying association between random objects that are located in general metric spaces. Special cases include functional and high-dimensional data as well as networks, covariance matrices and data on Riemannian manifolds, among other metric space-valued data. A key concept is the profile association, a measure based on distance profiles that intrinsically characterize the distributions of random objects in metric spaces. We rigorously establish a connection between the Hoeffding D statistic and the profile association and derive a permutation test with theoretical guarantees for consistency and power under alternatives to the null hypothesis of independence/no association. We extend this framework to the conditional setting, where the independence between random objects given a Euclidean predictor is of interest. In simulations across various metric spaces, the proposed profile independence test is found to outperform existing approaches. The practical utility of this framework is demonstrated with applications to brain connectivity networks derived from magnetic resonance imaging and age-at-death distributions for males and females obtained from human mortality data.
Keywords: Conditional association, Conditional independence test, Correlation, Distributional data, Functional data, High-dimensional data, Inference for networks, Non-Euclidean data, Permutation test
1 Introduction
Quantifying association and testing the independence between two random variables is a classical problem at the very core of statistics. The importance of this problem across many domains, including finance (Broock et al., 1996; Fan et al., 2013), genetics (Pritchard et al., 2000; Storey and Tibshirani, 2003; Browning and Browning, 2007) and many other areas is well known, and understanding the intricate relationships between variables is crucial for inference, prediction, and decision-making. Classical measures of dependence, such as Pearson’s correlation (Pearson, 1900), Spearman’s $\rho$ (Spearman, 1904) and Kendall’s $\tau$ (Kendall, 1938) have long served as fundamental tools for analyzing relationships between variables and are still very popular as they are simple and effectively capture linear or monotonic associations. Recently, the Chatterjee rank correlation coefficient (Chatterjee, 2021) has garnered significant attention and has been studied and applied across various fields (Bickel, 2022; Lin and Han, 2023; Han and Huang, 2024; Lin and Han, 2024).
As data collected in scientific research have become increasingly complex, reaching beyond traditional scalar or vector formats, there is a need for statistical methodology and especially inference that can be applied for data located in more general spaces (Dryden and Kent, 2015; Dubey and Müller, 2020; Marron and Dryden, 2021; Dubey et al., 2024). Notable examples for such non-traditional and non-Euclidean data are distribution-valued data and Riemannian manifold-valued data, which have numerous applications in fields such as economics (Gunsilius, 2023), biology (Bolstad et al., 2003), brain imaging (Lila et al., 2016; Lila and Aston, 2020) and geography (Mitchell et al., 2025). For data residing in spaces that are endowed with a (pseudo-)Riemannian geometry, a common approach is to apply transformations, such as the Riemannian log map, to project the data into a linear tangent space (Petersen and Müller, 2016; Bigot et al., 2017; Lin and Yao, 2019; Chen et al., 2023). If such a linearizing map is available, one may adopt traditional methodologies for data in Euclidean spaces such as quantifications for association of multivariate or functional data with their usual Euclidean (or $\mathcal{L}^{2}$) metrics. However, the linearizing maps that are commonly employed are either metric-distorting or non-bijective, requiring ad hoc solutions to construct pseudo-inverses (Bigot et al., 2017; Chen et al., 2023). Also data may reside in metric spaces without an underlying Riemannian geometry, such as the space of phylogenetic trees (Billera et al., 2001).
The study of data from general metric spaces has given rise to a prominent research area that has been referred to as metric statistics (Dubey et al., 2024), where random elements taking values in general metric spaces have been referred to as random objects (Müller, 2016; Bhattacharjee et al., 2025). The concept of a mean in general metric spaces was introduced by Fréchet (1948) and more recently extended to define conditional means in regression settings (Petersen and Müller, 2019) under the notion of Fréchet regression. These concepts are based on intrinsic metric functions and do not require any transformation to an extrinsic linear space. Such intrinsic approaches that directly confront the absence of linear operations have gained much interest in recent studies (Dubey and Müller, 2020; Zhu and Müller, 2023; Schötz, 2025; Jiang et al., 2024). Distance profiles (Dubey et al., 2024; Wang et al., 2024) corresponding to a class of one-dimensional distribution functions that characterize the distribution of random objects under mild conditions have proved instrumental for predictive conformal inference (Zhou and Müller, 2024a).
In this paper, we consider the setting where $X$ and $Y$ are random objects that may be situated in different metric spaces and develop a unified framework to quantify the association between $X$ and $Y$. We extend the concept of distance profiles to random object pairs $(X,Y)$ and demonstrate that the joint distance profile function equals the product of the marginal distance profiles of $X$ and $Y$ if and only if $X$ and $Y$ are independent. Building on this basic result, we introduce the profile association, a dependence measure that is zero if and only if $X$ and $Y$ are independent and is one when $X$ and $Y$ have a specific functional relationship that we characterize below. The analysis of the empirical profile association relies on a connection between the Hoeffding D statistic and the proposed profile association, which makes it possible to devise a consistent permutation test for testing the independence between $X$ and $Y$ and to derive its power properties, the profile independence test. Popular alternatives to obtain inference with distance-based tests include the energy test and the d-variable Hilbert-Schmidt Independence Criterion (dHSIC), which have been extended to non-Euclidean data (Székely et al., 2007; Lyons, 2013; Sejdinovic et al., 2013) and we include these methods in our numerical comparisons.
For random elements taking values in Banach spaces, Pan et al. (2020) introduced an independence measure based on the concept of ball covariance, a concept closely related to distance profiles. However, the proof of the independence-zero equivalence property of ball covariance in Pan et al. (2020) depends on the linear structure of the Banach space. In practical applications, it is uncommon to encounter data that reside exclusively in a Banach space without also belonging to a Hilbert space equipped with an inner product, where one has available an array of methods that utilize the availability of linear operations (Lee et al., 2020; Lundborg et al., 2022). We emphasize that our proposed approach to quantify and infer dependence also applies to high-dimensional and functional data equipped with the Euclidean or $L^{2}$ metric, as it is universally applicable across all metric spaces.
We note that recently Wang et al. (2024) extended the concept of ball covariance to metric spaces and this paper also features a subsection on its application for independence testing. However, since independence testing is only a secondary focus of this paper, the independence testing framework for random objects as presented remains incomplete, as the consistency properties along with the details and theoretical guarantees for the power of the resulting permutation implementation have not been established. Moreover, simulation studies demonstrate that our proposed profile independence test outperforms the ball covariance test across all settings. Very recently and independently of our work, another study on testing mutual independence between random objects was posted on arXiv (Chen and Dubey, 2025). While this work also utilizes distance profiles, it does not discuss the independence-zero equivalence property, and the proposed test statistic is based on estimated distance profiles. Our testing framework was derived simultaneously with this work, establishes the independence-zero equivalence, provides details for the dependence-1 equivalence, is based on a different Hoeffding D-statistic-based test statistic, employs different proof techniques and utilizes a half-permutation test for the study of consistency and power. Furthermore, we extend our approach to the important case of testing and quantifying conditional associations, when conditioning on a Euclidean random vector.
The quantification of conditional associations and inference is motivated by the fact that many real-world datasets involve complex random objects along with Euclidean predictors, such as age, time or experimental conditions. Understanding conditional relationships then enables researchers to model, predict, and interpret such data more effectively, especially in fields such as medical imaging, neuroscience, genetics, and environmental science. Specifically, we introduce the notion of conditional profile association to quantify the dependence between random objects $X$ and $Y$ given a Euclidean predictor $Z$ We showcase the application of conditional profile association for human longevity, where we study the relationship between the age-at-death distributions of males and females across different calendar years for various countries in Section 6.2. Using estimated joint and marginal distance profile functions, we also propose a statistic to carry out a conditional profile independence test to infer whether $X$ and $Y$ are conditionally independent and establish the uniform convergence of the estimated joint and marginal conditional distance profiles, which then implies the uniform convergence of conditional profile association. We also study the asymptotic distribution of the test statistic for the test for conditional independence under the null hypothesis and its implementation with a permutation test.
In summary, in this paper we introduce
  • A unified framework for testing and quantifying association between random objects with theoretical guarantees for the proposed profile association and profile independence test.
  • A half-permutation test procedure that is consistent under both the null and alternative hypotheses. We analyze its asymptotic power and demonstrate that it achieves a root-n minimax separation rate.
  • An extension to quantify and infer conditional association of random objects in the presence of Euclidean (scalar or vector) covariates. We establish uniform convergence rates for estimators for the conditional profile association, derive the asymptotic distribution of the conditional profile independence test statistic under the null hypothesis and study its implementation with a permutation test.
  • A demonstration of the practical merits of the proposed profile independence test and profile association measure through comprehensive simulation studies, demonstrating improved power performance compared to other existing methods, complemented by applications to brain connectivity network data and human mortality data. We establish the practical utility of the proposed conditional association measure and demonstrate that the proposed approach is not sensitive to the choice of the underlying metric.
The paper is organized as follows. In Section 2, we introduce the profile association for random objects along with some key properties. Section 3 presents the main methodology and theoretical results for empirical versions. We also discuss the proposed half permutation test for random objects and its asymptotic properties. In Section 4, we introduce test statistics and present asymptotic results for quantifying conditional independence in the presence of Euclidean covariates. In Section 5 we report the performance and comparison of the proposed methods for simulated data, followed by applications to brain imaging and human longevity that are presented in Section 6. Proofs and additional results are included in the Supplement.
2 Profile association for random objects
We state here some basic assumptions and notations. In the following, $X$ and $Y$ are random elements that take values in totally bounded, separable metric spaces $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with associated metrics $d_{1}(\cdot,\cdot)$ and $d_{2}(\cdot,\cdot)$, respectively. We refer to such random elements as random objects. We denote by $(S,\mathcal{S},\mathbb{P})$ an underlying probability space, where $\mathcal{S}$ is the Borel $\sigma$-algebra on $S$ and $\mathbb{P}$ is a probability measure, and by $\mu_{XY}$ the joint measure of $X$ and $Y$, where $\mu_{XY}(A\times B)=\mathbb{P}(s\in S:X(s)\in A\text{ and }Y(s)\in B)$ for any Borel measurable sets $A\subset\mathcal{M}_{1}$ and $B\subset\mathcal{M}_{2}$. We further assume that the marginal distributions of $X$ and $Y$, denoted by $\mu_{X}$ and $\mu_{Y}$, have finite first-order moments, i.e., $\int d_{1}(\omega_{1},x)\,\mathrm{d}\mu_{X}(x)<\infty$ and $\int d_{2}(\omega_{2},x)\,\mathrm{d}\mu_{Y}(x)<\infty$ for some (and thus any) $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
We are interested in quantifying the association between $X$ and $Y$, and especially in determining whether $X$ and $Y$ are independent, which is equivalent to $\mu_{XY}=\mu_{X}\times\mu_{Y}$. The following proposition states a necessary and sufficient condition for $X$ and $Y$ to be independent. In general metric spaces, linear operations like addition, subtraction and inner product that are fundamental to conventional statistical inference for Euclidean data are not available. Furthermore, in some metric spaces like $\mathcal{L}^{2}{[0,1]}$, the absence of a well-defined Lebesgue measure renders density-based inference methods infeasible. Since the only structure available in general metric spaces is the distance between two elements, developing intrinsic statistical methodologies that align with the underlying geometry requires leveraging these distances. The following proposition provides a sufficient condition under which the independence of probability measures is equivalent to the independence of distances between random elements and every point in the metric space.
Proposition 1.
If the probability measures $\mu_{X}$, $\mu_{Y}$, and $\mu_{XY}$ are fully determined by their values on closed balls, then $\mu_{XY}=\mu_{X}\times\mu_{Y}$ if and only if $d_{1}(X,\omega_{1})$ is independent of $d_{2}(Y,\omega_{2})$ for all $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
A probability measure $\mu$ is said to be “fully determined by its values on closed balls” if knowing the values $\mu(\overline{\mathcal{B}}(\omega,r))$ for all points $\omega\in\mathcal{M}$ and all radii $r>0$ is sufficient to uniquely specify $\mu$ on all Borel sets. This problem was first proposed and studied by Christensen (1970), Davies (1971), and Hoffmann-Jørgensen (1975), and has been recently investigated by Blanche et al. (2016), along with applications in statistics and probability discussed in Lyons (2013), Dubey et al. (2024), and Chen and Dubey (2025). The following proposition provides two sufficient conditions under which the condition in Proposition 1 holds. A probability measure $\mu$ on a metric space $(\mathcal{M},d)$ is said to be doubling if there exists a constant $L_{\mu}\geq 0$ such that
$\mu\left\{\overline{\mathcal{B}}(\omega,2r)\right\}\leq L_{\mu}\mu\left\{\overline{\mathcal{B}}(\omega,r)\right\}\text{ for all }r>0\text{ and }\omega\in\mathrm{supp}(\mu),$ (1)
where ¯$\overline{\mathcal{B}}(\omega,r):=\left\{\omega^{\prime}\in\mathcal{M}:d(\omega,\omega^{\prime})\leq r\right\}$ is the closed ball centered at $\omega$ with radius $r$, and $\mathrm{supp}(\mu):=\left\{\omega^{\prime}\in\mathcal{M}:\mu(\{\omega^{\prime}\})>0\right\}$ is the support of $\mu$. The concept of a doubling measure in metric spaces was first introduced by Beurling and Ahlfors (1956) and has since become a fundamental notion in analysis on metric spaces, including various areas such as harmonic analysis and geometric measure theory (Heinonen, 2001).
Proposition 2.
Proposition 1 holds provided that either of the following assumptions is satisfied:
  • (M0).
    $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are of strong negative type.
  • (P0).
    The measures $\mu_{X}$, $\mu_{Y}$ and $\mu_{XY}$ satisfy the doubling condition (1).
Assumption M0 pertains to the structural properties of the underlying metric spaces. The proof is based on Proposition 3.5, Proposition 3.7, and Lemma 3.8 in Lyons (2013). Preliminary background on metric spaces of strong negative type is provided in Appendix A, with additional details and illustrative examples available in Klebanov et al. (2005) and Lyons (2013). Assumption P0 concerns the properties of the associated probability measures and was employed in recent work Chen and Dubey (2025) to study mutual independence in metric spaces. Proposition 1 and Proposition 2 establish a connection between the independence of probability measures and the independence of distances of random elements, providing the basis for statistical inference for association and independence in metric spaces.
Our approach utilizes distance profiles (Dubey et al., 2024), where for each $\omega_{1}\in\mathcal{M}_{1}$, the marginal distance profile with respect to $\mu_{X}$ is defined as
$F^{X}_{\omega_{1}}(u)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u\right),\,\,\,u\geq 0,\,\,\omega_{1}\in\mathcal{M}_{1},$ (2)
and marginal distance profiles $F^{Y}_{\omega_{2}}(u)$ for each $\omega_{2}\in\mathcal{M}_{2}$ with respect to $\mu_{Y}$ are defined analogously. Distance profiles establish an intrinsic link between random objects and the one-dimensional distribution functions $F_{\omega}$ and provide a feasible framework for developing inference tools for random objects. This framework has been previously applied for the construction of two-sample tests (Dubey et al., 2024; Wang et al., 2024) and conformal prediction (Zhou and Müller, 2024a).
To measure the independence between $X$ and $Y$, we adopt the idea to extend (2) to the two-dimensional case by considering joint distance profiles for pairs of elements $(\omega_{1},\omega_{2})$,   $\omega_{1}\in\mathcal{M}_{1},\omega_{2}\in\mathcal{M}_{2}$, with respect to $\mu_{XY}$,
$F_{\omega_{1},\omega_{2}}^{XY}(u,v)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u,\,d_{1}(Y,\omega_{2})\leq v\right),\quad\,u,v\geq 0,\,\,\,\omega_{1}\in\mathcal{M}_{1},\,\,\omega_{2}\in\mathcal{M}_{2}.$ (3)
Proposition 2 then implies the following result, which relates the independence between random objects to their corresponding distance profiles.
Proposition 3.
Under either (M0) or (P0), the random objects $X$ and $Y$ are independent if and only if
$F_{\omega_{1},\omega_{2}}^{XY}(t,s)=F_{\omega_{1}}^{X}(t)F_{\omega_{2}}^{Y}(s),$
for all $t,s\geq 0$, $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
Proposition 3 enables us to transform the independence testing problem for random objects into testing a class of two-dimensional vectors $\{(d_{1}(X,\omega_{1}),d_{2}(Y,\omega_{2}))^{\top}\}$ indexed by $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$. Extensive research has been conducted on testing the independence of vector-valued random variables $X$ and $Y$ by quantifying the discrepancy between their joint measure and the product of their marginal measures. For scalar random variables $X$ and $Y$ with joint distribution function $F(x,y)$, in a seminal paper Hoeffding (1948) introduced the functional
$\Delta(F)=\iint\{F(x,y)-F(x,\infty)F(\infty,y)\}^{2}\,\mathrm{d}F(x,y),$
which corresponds to a type of profile association between $X$ and $Y$. When $F(x,y)$ has a continuous density function, $X$ and $Y$ are independent if and only if $\Delta(F)=0$.
When $X$ and $Y$ are random objects, one may analogously define their profile association as
$\Delta(X,Y):=\mathbb{E}_{X^{\prime},Y^{\prime}}\left[\iint\{F_{X^{\prime},Y^{\prime}}^{XY}(u,v)-F_{X^{\prime}}^{X}(u)F_{Y^{\prime}}^{Y}(v)\}^{2}{\rm d}F^{XY}_{X^{\prime},Y^{\prime}}(u,v)\right],$ (4)
where $(X^{\prime},Y^{\prime})$ is an independent copy of $(X,Y)$, and the random distance profiles $F_{X^{\prime},Y^{\prime}}^{XY}(u,v)$, $F_{X^{\prime}}^{X}(u)$, and $F_{Y^{\prime}}^{Y}(v)$ are defined as:
$F_{X^{\prime},Y^{\prime}}^{XY}(u,v):=\mathbb{P}_{X,Y}\left(d_{1}(X,X^{\prime})\leq u,\,d_{2}(Y,Y^{\prime})\leq v\right),$
$F_{X^{\prime}}^{X}(u):=\mathbb{P}_{X}\left(d_{1}(X,X^{\prime})\leq u\right),\quad F_{Y^{\prime}}^{Y}(v):=\mathbb{P}_{Y}\left(d_{2}(Y,Y^{\prime})\leq v\right),$
for all $u,v\geq 0$. Throughout this paper, the subscript in $\mathbb{E}_{X}$ (or $\mathbb{P}_{X}$) indicates that the expectation (or probability) is taken over the random variable(s) specified in the subscript only. For a random object $X$, i.e., a metric space-valued random variable, we denote by $X(s)\in\mathcal{M}$ the value it assumes in the metric space $\mathcal{M}$ for a specific argument $s$ that is an element of the underlying probability space $S$.
Theorem 1.
Assume that the density function $f^{XY}_{\omega_{1},\omega_{2}}(u,v)$ corresponding to the joint distribution function $F^{XY}_{\omega_{1},\omega_{2}}(u,v)$ exists and is continuous with respect to $\omega_{1}$, $\omega_{2}$, $u$, and $v$. Then
  • a).
    $\Delta(X,Y)=0$ if and only if the random variables $X$ and $Y$ are independent.
  • b).
    $\Delta(X,Y)$ attains its maximum value if and only if the real-valued random variables $d_{1}(\omega_{1},X)$ and $d_{2}(\omega_{2},Y)$ are comonotonic for all $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$, i.e.,
    $\{d_{1}(\omega_{1},X(s_{1}))-d_{1}(\omega_{1},X(s_{2}))\}\{d_{2}(\omega_{2},Y(s_{1}))-d_{2}(\omega_{2},Y(s_{2}))\}\geq 0,\text{ for all }s_{1},s_{2}\in S.$
    Furthermore, if $\mathcal{M}_{1}=\mathcal{M}_{2}$, $X$ and $Y$ are comonotonic if and only if $X=Y$ almost surely.
The assumption in Theorem 1 is relatively mild. Specifically, when $\mathcal{M}_{1}=\mathcal{M}_{2}=\mathbb{R}$, this assumption can be derived from the condition in Hoeffding (1948), which requires that the pair $(X,Y)$ has a continuous joint density function. The first statement of Theorem 1 shows that the proposed profile association measure satisfies both I-consistency and D-consistency; see Weihs et al. (2018):
  • I-consistency. If $X$ and $Y$ are independent, the profile association measure is zero.
  • D-consistency. If $X$ and $Y$ are dependent, the profile association measure is nonzero.
A similar independence-zero equivalence property has been established for ball covariance (Pan et al., 2020), which was proposed to test independence between Banach space-valued random elements (while we consider here random objects which take values in more general metric spaces). The proof of Theorem 1 relies on a mapping between negative-type metric spaces and Hilbert spaces (Schoenberg, 1937, 1938; Lyons, 2013).
Theorem 1 states that $\Delta(X,Y)$ achieves its maximum value if and only if $X=Y$ almost surely when $\mathcal{M}_{1}=\mathcal{M}_{2}$. Similar results have been established for other measures of dependence, including Hoeffding’s $D$ (Hoeffding, 1948), Bergsma–Dassios–Yanagimoto’s $\tau^{\ast}$ (Yanagimoto, 1970; Bergsma and Dassios, 2014; Weihs et al., 2018), and Chatterjee’s rank correlation (Chatterjee, 2021; Shi et al., 2022). These measures typically achieve their maximum value when $Y$ is a measurable or monotone function of $X$, rather than requiring $X$ and $Y$ to coincide almost surely.
The proposed profile association measure (4) specifically targets the quantification of association for random objects in general metric spaces. We note that monotone functions are not well-defined in metric spaces that lack a natural order and furthermore that alternative measures of independence often assign the same maximal association level to any monotone (or measurable) function of $X$, offering no additional insight into the degree of dependence between $Y$ and $X$. This limitation prevents these measures from distinguishing varying levels of dependence. In contrast, the proposed profile association uniquely identifies $Y=X$ almost surely as the maximal form of dependence, providing a more precise interpretation of maximal dependence. This property also indicates that test statistics based on (4) may have greater power in detecting dependence, and we find that this is indeed the case in the simulations reported in Section 5.
3 Methodology and theory
3.1 Testing independence for random objects
We are interested in testing the hypothesis
[…]
using random samples $\mathcal{D}_{\mathsf{XY}}:=\{(X_{i},Y_{i}):1\leq i\leq n\}$ drawn from the joint distribution of the random objects $(X,Y)$. The profile association $\Delta(X,Y)$ in (4) quantifies the strength of the relationship between random objects $X$ and $Y$ and provides the foundation for a test statistic to assess their independence.
Consider the functions $\psi_{1}:\mathcal{M}_{1}^{\otimes 4}\mapsto\{-1,0,1\}$ defined as
$\psi_{1}(x_{1},x_{2},x_{3},x_{4})=\mathds{1}_{\{d_{1}(x_{1},x_{3})\leq d_{1}(x_{1},x_{2})\}}-\mathds{1}_{\{d_{1}(x_{1},x_{4})\leq d_{1}(x_{1},x_{2})\}},\quad x_{i}\in\mathcal{M}_{1},$
and $\psi_{2}:\mathcal{M}_{2}^{\otimes 4}\mapsto\{-1,0,1\}$ defined as
$\psi_{2}(y_{1},y_{2},y_{3},y_{4})=\mathds{1}_{\{d_{2}(y_{1},y_{3})\leq d_{2}(y_{1},y_{2})\}}-\mathds{1}_{\{d_{2}(y_{1},y_{4})\leq d_{2}(y_{1},y_{2})\}},\quad y_{i}\in\mathcal{M}_{2}.$
For simplicity, we will omit the subscripts in $\psi_{1}$, $\psi_{2}$, $d_{1}$ and $d_{2}$, as the context will make it clear whether the functions are applied to elements in $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$. We also define the function
$h\big{(}(x_{1},y_{1}),\ldots,(x_{6},y_{6})\big{)}=$ $\frac{1}{4}\psi(x_{1},x_{2},x_{3},x_{4})\psi(x_{1},x_{2},x_{5},x_{6})$
×$\quad\times\psi(y_{1},y_{2},y_{3},y_{4})\psi(y_{1},y_{2},y_{5},y_{6}).$
In what follows, we use the notation $\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{6},Y_{6})\}$ to denote six independent copies of $(X,Y)$, where one has
$\mathbb{E}\Bigl{[}h\bigl{(}(X_{1},Y_{1}),\ldots,(X_{6},Y_{6})\bigr{)}\Bigr{]}\;=\;\Delta(X,Y).$
Given random samples $\mathcal{D}_{\mathsf{XY}}$, we employ a U-statistic as an estimator of $\Delta(X,Y)$, defined as
$D_{n}(\mathcal{D}_{\mathsf{XY}})=\frac{(n-6)!}{n!}\sum_{\mathcal{P}_{6}^{n}(\bm{\alpha})}h\left((X_{\alpha_{1}},Y_{\alpha_{1}}),\ldots,(X_{\alpha_{6}},Y_{\alpha_{6}})\right),$ (5)
where $\sum_{\mathcal{P}_{6}^{n}(\bm{\alpha})}$ demotes summation over all $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{6})^{\top}$ such that
$\alpha_{i}=1,\ldots,n;\,\,\alpha_{i}\neq\alpha_{j}\text{ if }i\neq j,\,\,\text{ for all }\,i,j=1,2,\ldots,6.$
We require some concepts related to the analysis of U-statistics, which were first introduced by Hoeffding (1948), with further details provided in Serfling (2009). Since the function $h$ is not symmetric, we first symmetrize it as follows,
$\mathfrak{h}((x_{1},y_{1}),\ldots,(x_{6},y_{6}))=\frac{1}{6!}\sum_{\mathcal{P}_{6}^{6}(\bm{\alpha})}h((x_{\alpha_{1}},y_{\alpha_{1}}),\ldots,(x_{\alpha_{6}},y_{\alpha_{6}})),$
where $\mathcal{P}_{6}^{6}(\bm{\alpha})$ denotes the set of all permutations of indices $\{1,2,\ldots,6\}$. For $r=1,2,\ldots,6$, let
$\mathfrak{h}_{r}((x_{1},y_{1}),\ldots,(x_{r},y_{r}))$
$=$ $\mathbb{E}\left[\mathfrak{h}((X_{1},Y_{1}),\ldots,(X_{6},Y_{6}))\mid X_{1}=x_{1},Y_{1}=y_{1},\ldots,X_{r}=x_{r},Y_{r}=y_{r}\right]$
and note that
$\mathbb{E}[\mathfrak{h}_{r}((X_{1},Y_{1}),\ldots,(X_{r},Y_{r}))]=\Delta(X,Y),\quad\text{for }r=1,2,\ldots,6.$
Define the centered version of $\mathfrak{h}_{r}$ as
$\tilde{\mathfrak{h}}_{r}((x_{1},y_{1}),\ldots,(x_{r},y_{r}))=\mathfrak{h}_{r}((x_{1},y_{1}),\ldots,(x_{r},y_{r}))-\Delta(X,Y),$
and
$\xi_{r}={\rm var}\big{[}\tilde{\mathfrak{h}}_{r}((X_{1},Y_{1}),\ldots,(X_{r},Y_{r}))\big{]},\quad r=1,2,\ldots,6.$
The variance terms $\xi_{r}$ play a key role in understanding the distributional properties of U-statistics, as they capture the variability of the centered kernels $\tilde{\mathfrak{h}}_{r}$. Specifically, $\xi_{1}$ corresponds to the variance of the first-order centered kernel and indicates whether the U-statistic has a non-degenerate linear term in its Hoeffding decomposition. When $\xi_{1}=0$ this is referred to as the degenerate case. This occurs when the first-order projection of the centered kernel function $\tilde{\mathfrak{h}}_{1}$ is zero, indicating that the statistic is orthogonal to all linear functions of the data. Then higher-order terms dominate the variance, often making the asymptotic distribution more complex. This situation often arises in independence testing, where the U-statistic is specifically designed to capture non-linear dependencies between random variables.
For $x_{1},x_{2}\in\mathcal{M}_{1}$, consider the functions $g_{1}:\mathcal{M}_{1}\times\mathcal{M}_{1}\to[0,1]$, defined as
$g_{1}(x_{1},x_{2}):=\mathbb{E}\left[\mathds{1}_{\{d(X,X^{\prime})\leq\max\{d(x_{1},X),d(x_{2},X^{\prime})\}\}}\right],$ (6)
where $X^{\prime}$ is an independent copy of $X$, and $\mathfrak{g}_{1}:\mathcal{M}_{1}\times\mathcal{M}_{1}\to\mathbb{R}$, defined as
$\mathfrak{g}_{1}(x_{1},x_{2}):=g_{1}(x_{1},x_{2})-\mathbb{E}g_{1}(x_{1},X)-\mathbb{E}g_{1}(X,x_{2})+\mathbb{E}g_{1}(X^{\prime},X).$ (7)
By substituting $x_{1},x_{2},X,X^{\prime}$ in (6) and (7) with $y_{1},y_{2},Y,Y^{\prime}$, one can similarly define $g_{2}(y_{1},y_{2})$ and $\mathfrak{g}_{2}(y_{1},y_{2})$, which map from $\mathcal{M}_{2}\times\mathcal{M}_{2}$ to $[0,1]$ and $\mathbb{R}$, respectively. The following result shows that the test statistic for the profile independence test is a degenerate U-statistic and establishes its asymptotic distribution under $\mathcal{H}_{0}$.
Theorem 2.
Under $\mathcal{H}_{0}$, the statistic $D_{n}(\mathcal{D}_{\mathsf{XY}})$, as defined in (5), is a degenerate U-statistic with $\xi_{1}=0$ and
$\xi_{2}=\frac{1}{225}\mathbb{E}[\mathfrak{g}^{2}_{1}(X_{1},X_{2})\mathfrak{g}^{2}_{2}(Y_{1},Y_{2})].$
Assuming that $\mathbb{E}[\mathfrak{g}^{2}_{1}(X_{1},X_{2})\mathfrak{g}^{2}_{2}(Y_{1},Y_{2})]\neq 0$, we have
[…]
where $Z_{j}$ are independent $\chi_{1}^{2}$ random variables, and $\lambda_{j}$ are the eigenvalues of the linear operator $\mathscr{H}:\mathcal{L}_{\mu_{XY}}^{2}(\mathcal{M}_{1}\otimes\mathcal{M}_{2})\to\mathcal{L}_{\mu_{XY}}^{2}(\mathcal{M}_{1}\otimes\mathcal{M}_{2})$, defined by
$\mathscr{H}(f)$ $=\iint\tilde{\mathfrak{h}}_{2}((x,y),(u,v))f(x,y){\rm d}\mu_{XY}(x,y),$
for all $(u,v)\in\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ and $f\in\mathcal{L}_{\mu_{XY}}^{2}(\mathcal{M}_{1}\otimes\mathcal{M}_{2})$.
3.2 Permutation test and asymptotic results
Since the asymptotic distribution in Theorem 2 is unknown, we use a permutation test to determine the critical values of the test statistic $D_{n}(\mathcal{D}_{\mathsf{XY}})$ when implementing the test. Permuting the data and recalculating the test statistic for each permutation is a standard approach to generate a reference distribution. In testing independence, the typical approach is to compute replicates of the test statistic by randomly permuting the $Y$ sample while keeping the order of the $X$ sample fixed (Székely et al., 2007). However, randomly permuting the entire $Y$ sample introduces dependence among the data pairs under the alternative $\mathcal{H}_{1}$ when $X$ and $Y$ are dependent, making it challenging to establish consistency of the critical value under $\mathcal{H}_{1}$. Here we adopt a half-permutation procedure introduced in Fan et al. (2024) that addresses this issue and does not require the splitting technique with its possible loss of power. Adopting this approach for our setting requires a different proof technique to obtain consistency and power bounds for the proposed profile independence test, as previous techniques are not applicable.
We use $\mathcal{P}_{\lfloor n/2\rfloor}^{n}$ to denote the set of permutations of $\{1,2,\ldots,n\}$ with cardinality $\lfloor n/2\rfloor$. Consider a sequence of i.i.d. random permutations $\Pi_{1},\ldots,\Pi_{B}$ that are independent of $(X,Y)$ and uniformly distributed on $\mathcal{P}_{\lfloor n/2\rfloor}^{n}$. We use $\Pi_{j}^{c}$ to denote the complement of $\Pi_{j}$ with respect to the set $\{1,2,\ldots,n\}$. Without loss of generality, we assume $n$ is even, so that $\Pi_{j}$ and $\Pi_{j}^{c}$ have the same cardinality $n/2$. For each $\Pi_{j}^{c}$, let $\{\sigma_{jk}\}_{k=1}^{N_{j}}$ be a sequence of random permutations on $\Pi_{j}^{c}$ that are independent of $(X,Y)$ and $\Pi_{j}$. Assume $\Pi_{j}=\{\imath_{1},\imath_{2},\ldots,\imath_{\lfloor n/2\rfloor}\}$ and $\sigma_{jk}\circ\Pi_{j}^{c}=\{\jmath_{1},\jmath_{2},\ldots,\jmath_{\lfloor n/2\rfloor}\}$, where $\sigma_{jk}\circ\Pi_{j}^{c}$ denotes the sequence of $\Pi_{j}^{c}$ under the permutation $\sigma_{jk}$. The permuted data corresponding to $(\Pi_{j},\sigma_{jk})$ are
$\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}):=\{(X_{\imath_{1}},Y_{\jmath_{1}}),(X_{\imath_{2}},Y_{\jmath_{2}}),\ldots,(X_{\imath_{\lfloor n/2\rfloor}},Y_{\jmath_{\lfloor n/2\rfloor}})\}.$
The test statistic based on permuted data $\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})$ is
$D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})):=\frac{(\lfloor n/2\rfloor-6)!}{\lfloor n/2\rfloor!}\sum_{\mathcal{P}_{6}^{\lfloor n/2\rfloor}(\bm{\alpha})}h((X_{\imath_{\alpha_{1}}},Y_{\jmath_{\alpha_{1}}}),\ldots,(X_{\imath_{\alpha_{6}}},Y_{\jmath_{\alpha_{6}}})).$
Let $N=\sum_{j=1}^{B}N_{j}$ be the total number of permutation pairs $(\Pi_{j},\sigma_{jk})$. Since the pairs $(\Pi_{j},\sigma_{jk})$ are random elements, we can define a randomized test at level $\alpha$ by
$\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}}):=\mathbb{P}\left(1+\sum_{(\Pi_{j},\sigma_{jk})}\mathds{1}_{\{n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})\leq\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}}\leq(1+N)\alpha\right),$ (8)
where the probability in (8) is taken over $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and the only randomness in $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ comes from the random permutations $(\Pi_{j},\sigma_{jk})$.
Under the null hypothesis, $X$ and $Y$ are independent, and the elements in the data sets $\mathcal{D}_{\mathsf{XY}}$ and $\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})$ are exchangeable. Consequently, $\{\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}_{j,k}$ and $n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})$ have the same asymptotic distribution by Theorem 2. Therefore, the probability on the right-hand side of equation (8) equals $\alpha$ for all distributions of $(X,Y)$ under the null hypothesis, which implies that $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ is a valid size-$\alpha$ test. For the consistency under the alternative $\mathcal{H}_{1}$, denote by $H(z)$ the cumulative distribution function of the asymptotic null distribution in Theorem 2. We employ the estimator
$\hat{H}_{N}(z)=\frac{1}{N}\sum_{(\Pi_{j},\sigma_{jk})}\mathds{1}_{\left\{\lfloor\frac{n}{2}\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\leq z\right\}}$ (9)
for $H(z)$ based on the permuted data, where $N=\sum_{j=1}^{B}N_{j}$ is the total number of permutation pairs $(\Pi_{j},\sigma_{jk})$. Theorem 3 below establishes the consistency of this estimator, which ensures that the size of the proposed profile independence test is correctly controlled under the alternative.
Theorem 3.
Under the conditions of Theorem 2 and if the number of permutations satisfies $N\rightarrow\infty$, it holds that
$|\hat{H}_{N}(z)-H(z)|=o_{P}(1)$
for every $z$ which is a continuity point of $H(z)$.
To further investigate the asymptotic power behavior of (8), we consider the alternative space $\mathcal{F}_{1}(\rho):=\{\mu_{XY}\in\mathcal{F}:\Delta(X,Y)\geq\rho\}$ for any positive $\rho$, where $\mathcal{F}$ denotes the subset of all probability measures on $\mathcal{M}_{1}\times\mathcal{M}_{2}$ whose cumulative distribution functions satisfy the assumptions in Theorem 1. Since we have already shown that the type-I error of ${\psi}_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ is always bounded by $\alpha$, we define the minimax separation radius as
$\rho^{\ast}(n,\alpha,\beta)=\inf\left\{\rho>0:\,\inf_{{\psi}\in\mathsf{\Psi}(\alpha)}\sup_{\mu_{XY}\in\mathcal{F}_{1}(\rho)}\mathbb{E}(1-\psi)\leq\beta\right\},$ (10)
where $\mathsf{\Psi}(\alpha)$ denotes the set containing all valid size-$\alpha$ independence tests and the expectation is taken over the joint probability measure $\mu_{XY}$. The separation (10) quantifies the smallest difference between the null and alternative hypotheses in terms of $\Delta(X,Y)$ for which the test can reliably distinguish between null and alternative under specified upper bounds on the type-I and type-II errors. The following result provides an upper bound for the separation distance $\rho^{\ast}(n,\alpha,\beta)$.
Theorem 4.
For any fixed $\alpha,\beta\in(0,1)$ such that $\alpha+\beta<1$, and assuming the number of permutations $N$ satisfies $N>(\alpha\beta)^{-1}-1$, there exists a constant $C$ that depends only on $\alpha$ and $\beta$, such that
$\rho^{\ast}(n,\alpha,\beta)\leq Cn^{-1/2}.$
4 Conditional independence test
Conditional independence testing is a statistical tool used to determine whether two variables are independent when conditioning on a third variable. It is used in a wide range of statistical applications where one studies complex relationships among variables, including causal inference (Zhang et al., 2011; Zhang et al., 2017; Peters et al., 2017), graphical models (Koller and Friedman, 2009; Williams and Mulder, 2020) and sufficient dimension reduction (Cook, 2004; Chen et al., 2015; Li, 2018). Many nonparametric conditional independence tests have been studied based on conditional density functions (Huang, 2010; Su and White, 2008; Banerjee et al., 2024), distribution functions (Linton and Gozalo, 1996; Zhou et al., 2020) and characteristic functions (Su and White, 2007; Wang et al., 2015; Wang and Hong, 2018). These methods are all restricted to the case where data reside in a vector space. To our knowledge, conditional independence tests for random elements in general metric spaces that do not support linear operations have remained unexplored.
In this section, we consider a setting where $X\in\mathcal{M}_{1}$ and $Y\in\mathcal{M}_{2}$ are random objects, while the conditional variable $Z$ is Euclidean, specifically $Z\in\mathcal{Z}$, where $\mathcal{Z}$ is a compact subset of ${\mathbb{R}}^{p}$. Our goal is to test the following hypothesis:
[…]
using i.i.d. samples $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{n}$ drawn from the joint distribution of $(X,Y,Z)$. To develop a suitable test statistic, we extend the concepts of marginal and joint distribution profiles to their conditional counterparts. For each $z\in\mathcal{Z}$, the marginal and joint conditional distribution profiles for $X$ and $Y$ are defined as
$F_{\omega_{1},z}^{X}(u)=\mathbb{P}\big{(}d(X,\omega_{1})\leq u\mid Z=z\big{)},\quad\text{ for all }\omega_{1}\in\mathcal{M}_{1},\,u\geq 0,$
$F_{\omega_{2},z}^{Y}(v)=\mathbb{P}\big{(}d(Y,\omega_{2})\leq v\mid Z=z\big{)},\quad\text{ for all }\omega_{2}\in\mathcal{M}_{2},\,v\geq 0;$
and
$F_{\omega_{1},\omega_{2},z}^{XY}(u,v)=$ $\mathbb{P}\big{(}d(X,\omega_{1})\leq u,\,d(Y,\omega_{2})\leq v\mid Z=z\big{)}$
for all $\omega_{1}\in\mathcal{M}_{1},\,\omega_{2}\in\mathcal{M}_{2},\,u,v\geq 0$.
Based on these conditional distance profiles, we define the conditional profile association
$\Delta_{XY}(z)=\mathbb{E}\left[\iint\left\{F_{X^{\prime},Y^{\prime},Z^{\prime}}^{XY}(u,v)-F_{X^{\prime},Z^{\prime}}^{X}(u)F_{Y^{\prime},Z^{\prime}}^{Y}(v)\right\}^{2}{\rm d}F_{X^{\prime},Y^{\prime},Z^{\prime}}^{XY}(u,v)\Bigm{|}Z^{\prime}=z\right],$ (12)
where $F_{X^{\prime},Y^{\prime},Z^{\prime}}^{XY}$, $F_{X^{\prime},Z^{\prime}}^{X}$ and $F_{Y^{\prime},Z^{\prime}}^{Y}$ are given by
$F_{X^{\prime},Y^{\prime},Z^{\prime}}^{XY}(u,v):=\mathbb{P}_{X,Y,Z}\left(d(X,X^{\prime})\leq u,\,d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right),$
$F_{X^{\prime},Z^{\prime}}^{X}(u):=\mathbb{P}_{X,Z}\left(d(X,X^{\prime})\leq u\mid Z=Z^{\prime}\right),\,F_{Y^{\prime},Z^{\prime}}^{Y}(v):=\mathbb{P}_{Y,Z}\left(d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right).$
We use $\Delta_{XY}(z)$ as measure of conditional profile association between $X$ and $Y$ given $Z=z$. When $Z$ is a continuous random variable, such as time, conditional profile association can be used to quantify the time-varying behavior of the dependency between $X$ and $Y$ as a function of $Z$. Such measures have important applications in many areas, including the life and social sciences and climatology. We provide a specific application example for the analysis of human longevity in Section 6.
We focus here on the case where $Z$ is univariate and continuous. To avoid the curse of dimensionality under some structural assumptions, for the multivariate case one can apply single index Fréchet regression (Bhattacharjee and Müller, 2023), employing arguments that are similar to those in Zhou and Müller (2024a) for the case of multivariate predictors in conformal inference. For the univariate case we use nonparametric regression with argument $Z$ to implement conditional profiles, employing local linear estimators (Fan and Gijbels, 1992) for $F_{\omega_{1},z}^{X}(u)$, $F_{\omega_{2},z}^{Y}(v)$, and $F_{\omega_{1},\omega_{2},z}^{XY}(u,v)$. This yields
$\hat{F}^{X}_{\omega_{1},z}(u)=$ $\mathrel{\mathop{\arg\min}\limits_{\beta_{0}\in\mathbb{R}}}\frac{1}{nh_{n}}\sum_{i=1}^{n}\left\{\mathds{1}_{\{d(X_{i},\omega_{1})\leq u\}}-\beta_{0}-\beta_{1}(Z_{i}-z)\right\}^{2}\mathrm{K}\left(\frac{Z_{i}-z}{h_{n}}\right);$
$\hat{F}^{Y}_{\omega_{2},z}(v)=$ $\mathrel{\mathop{\arg\min}\limits_{\beta_{0}\in\mathbb{R}}}\frac{1}{nh_{n}}\sum_{i=1}^{n}\left\{\mathds{1}_{\{d(Y_{i},\omega_{2})\leq v\}}-\beta_{0}-\beta_{1}(Z_{i}-z)\right\}^{2}\mathrm{K}\left(\frac{Z_{i}-z}{h_{n}}\right);$
$\hat{F}^{XY}_{\omega_{1},\omega_{2},z}(u,v)=$ $\mathrel{\mathop{\arg\min}\limits_{\beta_{0}\in\mathbb{R}}}\frac{1}{nh_{n}}\sum_{i=1}^{n}\left\{\mathds{1}_{\{d(X_{i},\omega_{1})\leq u\}}\mathds{1}_{\{d(Y_{i},\omega_{2})\leq v\}}-\beta_{0}-\beta_{1}(Z_{i}-z)\right\}^{2}$
×$\quad\quad\quad\quad\times\mathrm{K}\left(\frac{Z_{i}-z}{h_{n}}\right),\hskip 369.88582pt$
where $\mathrm{K}(\cdot)$ is a symmetric and Lipschitz continuous density kernel on $[-1,1]$ and $h_{n}$ is a sequence of bandwidths.
To study the asymptotic behavior of these conditional distance profile estimators, we introduce the sets of functions
$\mathcal{F}_{1}$ =$=\,\{f(x)=\mathds{1}_{\{d(\omega_{1},x)\leq t\}}:\omega_{1}\in\mathcal{M}_{1}\,\,\,t\in{\mathbb{R}}^{+}\},$
$\mathcal{F}_{2}$ =$=\,\{f(y)=\mathds{1}_{\{d(\omega_{2},y)\leq t\}}:\omega_{2}\in\mathcal{M}_{2},\,\,\,t\in{\mathbb{R}}^{+}\},$
which are classes of indicator functions indexed by $\omega_{1}$, $\omega_{2}$ and $t$. For a bounded function class $\mathcal{F}$ defined on $\mathcal{M}$, the uniform covering number is $\mathcal{N}(\epsilon,\mathcal{F}):=\sup_{\mathbb{Q}}N(\epsilon,\mathcal{F},d_{\mathbb{Q}}),$ where the supremum is taken over all probability measures $\mathbb{Q}$ on $\mathcal{M}$ with bounded second-order moments. Here, $d_{\mathbb{Q}}$ is the $\mathcal{L}_{\mathbb{Q}}^{2}$ metric, defined for any two functions $f,g\in\mathcal{F}$ as: $d_{\mathbb{Q}}^{2}(f,g)=\int\{f(x)-g(x)\}^{2}\,\mathrm{d}\mathbb{Q}(x).$ We require the following regularity assumptions.
Assumption 1.
The marginal distribution of $Z$ has a continuous density function $f_{Z}$, which satisfies $\inf_{z\in\operatorname{Support}(f_{Z})}f_{Z}(z)>c_{1}$ and $\sup_{z\in\mathcal{X}}f_{Z}(z)<c_{2}$ for strict positive constants $c_{1}$ and $c_{2}$.
Assumption 2.
The bandwidth sequence $\{h_{n}\}_{n\geq 1}$ satisfies $nh_{n}/\log n\rightarrow\infty$ and $|\log h_{n}|/\log\log n\rightarrow\infty$ as $n\rightarrow\infty$.
Assumption 3.
For every $\omega_{1}\in\mathcal{M}$, $\omega_{2}\in\mathcal{M}_{2}$ and $z\in\mathcal{X}$, the distance profile $F^{XY}_{\omega_{1},\omega_{2},z}$ is absolutely continuous with continuous density $f_{\omega_{1},\omega_{2},z}$ and there exist strict positive constants $c_{3}$ and $c_{4}$ such that $\inf_{t\in\operatorname{support}\left(f_{\omega_{1},\omega_{2},z}\right)}f_{\omega_{1},\omega_{2},z}(t)\geq c_{3}$ and $\sup_{t\in\mathbb{R}^{+}}f_{\omega_{1},\omega_{2},z}(t)\leq c_{4}<\infty$.
Assumption 4.
For every $\omega_{1}\in\mathcal{M}_{1}$, $\omega_{2}\in\mathcal{M}_{2}$ and $u,v\in{\mathbb{R}}^{+}$, $F^{XY}_{\omega_{1},\omega_{2},z}(u,v)$ is second order differentiable and has bounded second order derivatives with respect to $z$. Moreover, there exists a constant $L^{\prime}$ such that $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1}^{\prime},\omega_{2},z}(u,v)|\leq L^{\prime}d(\omega_{1},\omega_{1}^{\prime})$, $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1},\omega_{2}^{\prime},z}(u,v)|\leq L^{\prime}d(\omega_{2},\omega_{2}^{\prime})$ for all $z\in\mathcal{X},u,v\in{\mathbb{R}}^{+}$ and $\omega_{1},\omega_{1}^{\prime},\in\mathcal{M}_{1}$, $\omega_{2},\omega_{2}^{\prime},\in\mathcal{M}_{2}$.
Proposition 4.
Under assumptions 1 and 2, if $\max\{\mathcal{N}(\epsilon,\mathcal{F}_{1}),\,\mathcal{N}(\epsilon,\mathcal{F}_{2})\}\lesssim\epsilon^{-v}$ for some $v>0$ and $h_{n}\asymp(n/\log n)^{-1/5}$, then:
[…]
The first two statements of Proposition 4 follow directly from Theorem 3 in Zhou and Müller (2024a). Under the assumption $\max\{\mathcal{N}(\epsilon,\mathcal{F}_{1}),\,\mathcal{N}(\epsilon,\mathcal{F}_{2})\}\lesssim\epsilon^{-v}$, Lemma A.1 in Einmahl and Mason (2000) and Lemma S.4 in Zhou and Müller (2024b) imply that the function class $\mathcal{F}_{12}=\left\{f(x)=\mathds{1}_{\{d(\omega_{1},x)\leq t_{1}\}}\mathds{1}_{\{d(\omega_{2},x)\leq t_{2}\}}:\omega_{1}\in\mathcal{M}_{1},\,\omega_{2}\in\mathcal{M}_{1},\,t_{1},t_{2}\in\mathbb{R}\right\}$ has a polynomial covering number. The third statement is a consequence of Lemma 1 and Theorem 3 in Zhou and Müller (2024a).
An estimator for the conditional profile association ${\Delta}_{XY}(z)$ is then obtained through another nonparametric regression step,
$\hat{\Delta}_{XY}(z)=\mathrel{\mathop{\arg\min}\limits_{\beta_{0}\in\mathbb{R}}}\frac{1}{nh_{n}}\sum_{j=1}^{n}\left\{\hat{R}_{j}-\beta_{0}-\beta_{1}(Z_{j}-z)\right\}^{2}\mathrm{K}\left(\frac{Z_{j}-z}{h_{n}}\right),$ (13)
where
$\hat{R}_{j}=\iint\left\{\hat{F}_{X_{j},Y_{j},Z_{j}}^{XY}(u,v)-\hat{F}_{X_{j},Z_{j}}^{X}(u)\hat{F}_{Y_{j},Z_{j}}^{Y}(v)\right\}^{2}{\rm d}u{\rm d}v.$
The following result provides the uniform convergence of this estimator.
Theorem 5.
Under the assumptions of Proposition 4,
[…]
This result provides theoretical justification for the estimator (13) and supports its use as a measure for the conditional association of random effects. Using the estimated conditional profiles we then obtain the following test statistic for testing (11),
$\mathcal{T}_{n}=\frac{1}{n}\sum_{i=1}^{n}\iint\left\{\hat{F}^{XY}_{X_{i},Y_{i},Z_{i}}(u,v)-\hat{F}^{X}_{X_{i},Z_{i}}(u)\hat{F}^{Y}_{Y_{i},Z_{i}}(v)\right\}^{2}{\rm d}\hat{F}_{X_{i},Y_{i},Z_{i}}^{XY}(u,v).$ (14)
To study the asymptotic distribution of $\mathcal{T}_{n}$, we define the following covariance operator
$\mathcal{C}_{x,y,z}(u_{1},u_{2},v_{1},v_{2})\,=\,$ $C^{XY}_{x,y,z}(u_{1},u_{2},v_{1},v_{2})+F^{X}_{x,z}(u_{1})F^{X}_{x,z}(u_{2})C^{Y}_{y,z}(v_{1},v_{2})$ (15)
$+F^{Y}_{y,z}(v_{1})F^{Y}_{y,z}(v_{2})C^{X}_{x,z}(u_{1},v_{2})+\mathcal{C}_{1}(u_{1},u_{2},v_{1},v_{2})+\mathcal{C}_{2}(u_{1},u_{2},v_{1},v_{2}),$
where
$C^{XY}_{x,y,z}(u_{1},u_{2},v_{1},v_{2})$ $=\mathrm{Cov}\left(\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(x,X)\leq u_{1}\}}\mathds{1}_{\{d(y,Y)\leq v_{1}\}},\right.$
$\qquad\qquad\left.\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(x,X)\leq u_{2}\}}\mathds{1}_{\{d(y,Y)\leq v_{2}\}}\right);$
$C^{X}_{x,z}(u_{1},u_{2})$ $=\mathrm{Cov}\left(\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(x,X)\leq u_{1}\}},\right.$
$\qquad\qquad\left.\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(x,X)\leq u_{2}\}}\right);$
$C^{Y}_{y,z}(v_{1},v_{2})$ $=\mathrm{Cov}\left(\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(y,Y)\leq v_{1}\}},\right.$
$\qquad\qquad\left.\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(y,Y)\leq v_{2}\}}\right).$
$\mathcal{C}_{1}(u_{1},u_{2},v_{1},v_{2})$ $=\mathrm{Cov}\left(\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(x,X)\leq u_{1}\}}\mathds{1}_{\{d(y,Y)\leq v_{1}\}},\right.$
$\qquad\qquad\left.F_{x,z}^{X}(u_{2})\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(y,Y)\leq v_{2}\}}\right)$
$\mathcal{C}_{2}(u_{1},u_{2},v_{1},v_{2})$ $=\mathrm{Cov}\left(\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(x,X)\leq u_{1}\}}\mathds{1}_{\{d(y,Y)\leq v_{1}\}},\right.$
$\qquad\qquad\left.F_{y,z}^{Y}(v_{2})\frac{1}{f_{Z}(z)}\mathrm{K}\left(\frac{Z-z}{h}\right)\mathds{1}_{\{d(x,X)\leq u_{2}\}}\right).$
Theorem 6.
Under Assumptions 1 to 4, $nh^{2}\log n\rightarrow 0$ and $\mathcal{H}_{0,c}$, $nh\mathcal{T}_{n}$ converges in distribution to the law of a random variable $L=\sum_{k=1}^{\infty}Z_{k}^{2}\mathbb{E}_{X,Y,Z}(\lambda_{k}^{X,Y,Z})$, where $Z_{1},Z_{2},\ldots$ is a sequence of i.i.d. $\mathcal{N}(0,1)$ random variables, and $\lambda_{1}^{x,y,z}\geq\lambda_{2}^{x,y,z}\geq\cdots$ are the eigenvalues of the covariance operator (15).
Since the limit distribution in the theorem above is unknown, we implement the test using a permutation approach as introduced in Section 3.2. The permuted data are
$\mathcal{D}_{\mathsf{XYZ}}(\Pi_{j},\sigma_{jk}):=\{(X_{\imath_{1}},Y_{\jmath_{1}},Z_{\jmath_{1}}),(X_{\imath_{2}},Y_{\jmath_{2}},Z_{\jmath_{2}}),\ldots,(X_{\imath_{\lfloor n/2\rfloor}},Y_{\jmath_{\lfloor n/2\rfloor}},Z_{\jmath_{\lfloor n/2\rfloor}})\}.$
This permutation procedure is similar to that presented in Berrett et al. (2020). Let $\mathcal{T}_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XYZ}}(\Pi_{j},\sigma_{jk}))$ represent the test statistic defined in (14) based on the permuted data $\mathcal{D}_{\mathsf{XYZ}}(\Pi_{j},\sigma_{jk})$, and let $\mathcal{T}_{n}(\mathcal{D}_{\mathsf{XYZ}})$ be the corresponding test statistic calculated using the original data. The $p$-value resulting from the test is then
$p_{c}:=\frac{1+\sum_{\Pi_{j},\sigma_{jk}}\mathds{1}_{\{\lfloor n/2\rfloor\mathcal{T}_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XYZ}}(\Pi_{j},\sigma_{jk}))\geq n\mathcal{T}_{n}(\mathcal{D}_{\mathsf{XYZ}})\}}}{1+N},$
where $N$ denotes the total number of permutations. The randomized test for testing (11) at significance level $\alpha$ is then defined as $\psi_{\alpha}^{c}:=\mathbb{P}(p_{c}\leq\alpha)$.
Denoting by $\Gamma(z)$ the cumulative distribution function the limiting distribution in Theorem 6, its corresponding permutation-based estimator is
$\hat{\Gamma}_{N}(z)=\frac{1}{N}\sum_{(\Pi_{j},\sigma_{jk})}\mathds{1}_{\left\{\frac{n}{2}\mathcal{T}_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XYZ}}(\Pi_{j},\sigma_{jk}))\leq z\right\}},$
for which we obtain the following convergence result.
Corollary 1.
Under the same assumptions as in Theorem 6, $|\hat{\Gamma}_{N}(z)-\Gamma(z)|=o_{P}(1)$ holds for every continuity point $z$ of $\Gamma(z)$, and $\mathbb{P}(\psi_{\alpha}^{c})\leq\alpha$.
5 Simulations
We report here finite sample comparisons of the performance of the proposed profile approach with various alternative methods when $X$ and $Y$ are located in pertinent metric spaces. These include the Euclidean space $\mathbb{R}^{p}$, the sphere $\mathbb{S}^{2}$ and the 2-Wasserstein space of one-dimensional distributions $\mathbb{W}^{2}$ with the Wasserstein metric, where the latter is
$d_{W}(x,y)^{2}=\int_{0}^{1}\{Q_{x}(u)-Q_{y}(u)\}^{2}\,{\rm d}u$ (16)
and $Q_{x},\,Q_{y}$ are the quantile functions associated with the distributions $x$ and $y$.
For each setting, both the proposed and comparison methods were implemented using the permutation test described in Section 3.2 and empirical power curves at significance level 0.05 are obtained from 500 Monte Carlo runs. We consider the following settings, where in all settings the dependence is quantified by the parameter $\rho$, where $\rho=0$ corresponds to independence and the level of dependence (association) increases with $\rho$. We obtained empirical power curves that illustrate how the power of the profile independence test increases as the argument $\rho$ increases, where we consider $0\leq\rho\leq 1$
  • $\mathbb{R}^{1}$-Cir: Denoting by $\|v\|_{\zeta}$ the $\zeta$-norm for any Euclidean vector $v$. For any $1\leq\zeta\leq\infty$ we consider pairs $(X,Y)$ of one-dimensional random variables $X$ and $Y$ that satisfy the condition $\|(X,Y)\|_{2/\rho}\leq 1$, where $X$ and $Y$ are independently drawn from $\text{Unif}(-1,1)$. When $\rho=0$, $(X,Y)$ is uniformly distributed over the square $[-1,1]\times[-1,1]$. When $\rho=1$, $(X,Y)$ follows a uniform distribution within the unit circle $\{(x,y)\mid x^{2}+y^{2}\leq 1\}$, with domains for the uniform distribution located in between the square and the circle for $10<\rho<$ 1 .
  • $\mathbb{R}^{1}$-Bimodal: The pair $(X,Y)$ of one-dimensional random variables satisfies $Y=\epsilon|X|^{\rho}$, where $X$ and $\epsilon$ are independent and follow mixture distributions $0.2\mathcal{N}(-3,1)+0.8\mathcal{N}(3,1)$.
  • $\mathbb{R}^{p}$-Log: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are related only through their first elements, specifically $Y_{1}=\rho\log(4X_{1}^{2})+\epsilon$. The variables $Y_{2},\ldots,Y_{p}$, $X_{1},X_{2},\ldots,X_{p}$, and $\epsilon$ are all distributed as $\mathcal{N}(0,1)$ and are independent.
  • $\mathbb{R}^{p}$-Pow: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are related only through the first element, where $Y_{1}=\epsilon\times(|X_{1}|^{\rho})$. The variables $Y_{2},\ldots,Y_{p}$ and $\epsilon$ are distributed as $\mathcal{N}(0,1)$, while $X_{1},X_{2},\ldots,X_{p}$ are distributed as $\mathrm{Unif}(-1,1)$. Additionally, $Y_{2},\ldots,Y_{p}$, $X_{1},X_{2},\ldots,X_{p}$, and $\epsilon$ are independent.
  • $\mathbb{S}^{2}$-Lin: $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by independent standard normal random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$:
    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
    $Y_{1}=\frac{\rho\epsilon_{1}+(1-\rho)\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|},$
    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(\rho\epsilon_{1}+(1-\rho)\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
  • $\mathbb{S}^{2}$-Log: $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by independent standard normal random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$:
    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
    $Y_{1}=\frac{\rho\log(4\epsilon_{1}^{2})+\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|},$
    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(\rho\log(4\epsilon_{1}^{2})+\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
  • $\mathbb{S}^{2}$-Pow: $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by independent standard normal random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$:
    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
    $Y_{1}=\frac{|\epsilon_{1}|^{\rho}+\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|},$
    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(|\epsilon_{1}|^{\rho}+\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
  • $\mathbb{S}^{2}$-Pow-Bimodal: Here $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by the random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$, where $\epsilon_{i}$ are random samples from a mixture of Gaussians $0.2\mathcal{N}(-1,1)+0.8\mathcal{N}(1,1)$:
    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
    $Y_{1}=\frac{|\epsilon_{1}|^{\rho}+\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|}$
    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(|\epsilon_{1}|^{\rho}+\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
  • $\mathbb{W}^{2}$-Mean: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(\mu_{X},1)$ and $Y=\mathcal{N}(\mu_{Y},1)$, where $\mu_{X}\sim\mathrm{Unif}(-1,1)$, $\mu_{Y}=\epsilon\times|\mu_{X}|^{\rho}$, and $\epsilon\sim\mathcal{N}(0,1)$ and is independent of $\mu_{X}$.
  • $\mathbb{W}^{2}$-Var: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(0,\sigma_{X}^{2})$ and $Y=\mathcal{N}(0,\sigma_{Y}^{2})$, where $\sigma_{X}^{2}\sim\mathrm{Unif}(0,1)$ and $\sigma_{Y}=\rho\sin(\sigma_{X}\pi)+\epsilon$, and where $\epsilon\sim\mathrm{Unif}(0,2)$ and is independent of $\sigma_{X}$.
For cases where $X$ and $Y$ are in $\mathbb{R}^{p}$ for $p\geq 1$, we compare the proposed profile independence test with the energy test (Székely et al., 2007), the ball divergence test (Pan et al., 2020) and Chatterjee’s rank test (Chatterjee, 2021). All methods are implemented using R packages: the energy test (Energy) via dcov.test() from the energy package, the ball divergence test (Ball) via bcov.test() through the Ball package (Zhu et al., 2021) and Chatterjee’s rank test (Chatterjee) via xicor() through the XICOR package. To compare different approaches, we focus on the power curves as function of $\rho$ for the simulation settings described above. Figure 1 indicates that the proposed method outperforms all comparable methods, particularly in the $\mathbb{R}^{1}$-Bimodal setting, where the energy test and Chatterjee’s rank test fail. Additionally, in this more complex setting, the proposed test exhibits a faster increase in power as the sample size increases.
Figure 2 shows that the proposed profile independence test (blue) continues to consistently outperform the energy test (black) and the ball divergence test (red) across all simulation settings where $X$ and $Y$ reside in multivariate Euclidean spaces or metric spaces. The simulations thus provide evidence that the profile independence test effectively adapts to complex dependency structures where other methods fail. For example, when the relationship between $X$ and $Y$ is linear, a scenario where also classical methods perform well, the energy test performs comparably to the profile independence test when the relationship between $X$ and $Y$ is linear. However, in more complex cases, such as log- or power-based dependencies, and especially in settings $\mathbb{S}^{2}$-Log and $\mathbb{S}^{2}$-Pow, the energy test performs poorly, whereas the profile independence test maintains strong detection power. These results provide evidence for the robustness and efficiency of the proposed test in detecting dependencies across diverse scenarios.
Figure 1: Power curves comparing the proposed profile independence test (blue), the ball statistic test (red), the Chatterjee est (yellow) and the energy test (black), where power is a function of the increasing dependency parameter $\rho$, for two simulation settings: $\mathbb{R}^{1}$-Cir (first row) and $\mathbb{R}^{1}$-Bimodal (second row) The dashed grey line indicates the target significance level of $0.05$. Three different sample sizes are considered: $n=100$ (left column), $n=200$ (middle column), and $n=400$ (right column).
Figure 2: Power curves comparing the proposed profile independence test (blue), the ball statistic test (red), the Chatterjee test (yellow) and the energy test (black), where power is a function of the increasing dependency parameter $\rho$, for nine simulation settings: $\mathbb{R}^{5}$-Log, $\mathbb{R}^{10}$-Log, and $\mathbb{R}^{10}$-Pow (left, middle, right panel in first row); $\mathbb{S}^{2}$-Lin, $\mathbb{S}^{2}$-Log, and $\mathbb{S}^{2}$-Pow (left, middle, right panel in second row); $\mathbb{S}^{2}$-Pow-Bimodal, $\mathbb{W}^{2}$-Mean, and $\mathbb{W}^{2}$-Var (left, middle, right panel in third row). The dashed grey line shows the target significance level of $0.05$, and the sample size is set to $n=400$.
6 Data illustrations
6.1 Connectivity networks in fMRI based brain imaging
Functional MRI (fMRI) imaging data acquisition is a well-established and important tool for the study of brain function, due to its high temporal resolution. It is deployed to capture spontaneous neural activity during the resting state and also for task-specific responses to external stimuli. In general, fMRI produces complex data that require extensive pre-processing and then can be viewed as connectivity matrices, random densities or signals that reflect intricate spatial and temporal dynamics. Here we apply the proposed profile independence test to investigate the relationship between resting-state and task-evoked fMRI brain connectivity using data from the HCP 1200 subject release (Van Essen et al., 2013; Elam et al., 2021). Understanding this association reveals how the brain’s functional architecture varies across different tasks compared to its intrinsic organization that is thought to be reflected during the resting state.
The data are publicly accessible at https://www.humanconnectome.org. For our analysis, we focus on a subset comprising $n=209$ young, healthy adults who completed both resting-state and motor task fMRI sessions. The resting-state fMRI data were collected while participants were instructed to keep their eyes open and fixate on a projected bright cross-hair on a dark background in a darkened room. For the motor task fMRI recordings, participants were presented with visual cues prompting them to perform specific movements, such as squeezing their left or right toes or moving their tongue, to map specific motor areas. Each time block for the recording of a specific movement type lasted 12 seconds, included 10 movements and was preceded by a 3-second cue. After preprocessing, the average blood-oxygen-level-dependent (BOLD) signals were recorded across different voxels at 284 time points. Further details about the experiment, preprocessing and data acquisition are available in the reference manual for the WU-Minn HCP 1200 Subjects Data Release (Wu-Minn, 2017).
Our analysis focuses on nine regions of interest, nucleus accumbens, amygdala, caudate, cerebellum, diencephalon, hippocampus, pallidum, putamen, and thalamus, in both left and right brain hemispheres. Correlation matrices of BOLD signals were constructed as per established practice in fMRI analysis, namely calculating the normalized cross-correlations of fMRI time series based on the voxels residing around a seed region for each region of interest, separately for left and right brain hemispheres. For the task-evoked fMRI signals, the correlation matrices were calculated using the time segments corresponding to each specific task. For instance, the correlation matrix for “squeezing left toe” was derived from the time periods during which participants were actively performing the “squeezing left toe” task. For each subject, this resulted in a total of eight $9\times 9$ correlation matrices, corresponding to one rest phase and three task phases (resting state, squeezing left toe, squeezing right toe, and moving tongue) for each of the two brain hemispheres (left and right).
Figure 3 shows the estimated Fréchet mean of these eight covariance matrices when employing the Frobenius norm. This indicates that the task-evoked correlation matrices display higher correlations between different brain regions compared with the resting-state correlation matrices.
Figure 3: Fréchet mean correlation matrices for resting-state fMRI (left-most column) and task-evoked fMRI (next three columns: squeezing left toe, squeezing right toe, and moving tongue). The top row represents the left hemisphere and the bottom row the right hemisphere.
We then proceeded to apply the proposed profile independence test, along with the energy test (Székely et al., 2007), aiming to test the independence of random object pairs $\{(X_{i},Y_{i})\}_{i=1}^{209}$. Here $X_{i}$ represents the correlation matrix corresponding to resting-state fMRI, and $Y_{i}$ the correlation matrix corresponding to task-evoked fMRI, for the various motor tasks for the $i$th subject. We consider three metrics for the space of $9\times 9$ positive definite matrices $A$ and $B$. Specifically, we compare the results for the following three metrics that were applied previously in the literature for various data analysis tasks:
  • a)
    Frobenius metric
    $d_{F}({A},{B})=\|A-B\|_{F}=\sqrt{\operatorname{tr}\left[({A}-{B})({A}-{B})^{\top}\right]}.$
  • b)
    Affine Invariant Riemannian Metric (AIRM) (Moakher, 2005)
    $d_{R}(A,B)=\|\log(A^{-1}B)\|_{F}.$
  • c)
    Log-Cholesky metric (Lin, 2019)
    $d_{C}(A,B)=\left\{\left\|\lfloor L_{A}\rfloor-\lfloor L_{B}\rfloor\right\|_{F}^{2}+\sum_{j=1}^{9}(\log L_{A,jj}-\log L_{B,jj})^{2}\right\}^{1/2},$
    where $L_{A}$ and $L_{B}$ are the Cholesky decompositions of $A$ and $B$, respectively. Here $\lfloor L_{A}\rfloor$ denotes the off-diagonal part of $L_{A}$ and $L_{A,jj}$ the $j$th diagonal element of $L_{A}$.
For each metric and method, we conducted six tests in total: three for each motor task (“squeezing the left toe”, “squeezing the right toe”, and “moving the tongue”) separately for the left and right hemispheres. The $p$-values resulting from the permutation test implementation of the proposed profile independence test are shown in Table 1 and are also compared with the results from the energy test for independence. For the tongue movement task (columns 3, 6, and 9), both the proposed profile independence test and the energy test yield the same result across all three metrics: there is no evidence to reject the null hypothesis that the correlation matrices between the resting state and the tongue movement task are independent. For the toe-squeezing task, under the Frobenius metric (columns 1 and 2), both the proposed profile independence test and the energy test indicate dependence at the 0.05 significance level. However, the energy test fails to detect dependence under the AIRM and Log-Cholesky metrics (row 2, columns 5 and 8), whereas the proposed profile independence test consistently rejects the null hypothesis in all these cases. These results demonstrate the consistency of the proposed test across different metric choices.
Table 1: Comparison of $p$-values obtained when applying the proposed profile independence test and the energy test for independence for the null hypothesis of independence between resting-state fMRI and task-evoked fMRI connectivity matrices for three motor tasks: $\mathrm{Ltoe}$ (squeezing the left toe), $\mathrm{Rtoe}$ (squeezing the right toe), and $\mathrm{tongue}$ (moving the tongue). The results are presented for left and right hemispheres and for three different metrics: Frobenius metric (left three columns), Affine-invariant Riemannian metric (middle three columns) and Log-Cholesky metric (right three columns). The first two rows correspond to correlation matrices show the results for the left hemisphere and the last two rows for the right hemisphere. The $p$-values for the proposed profile independence test and the energy test were obtained through permutation tests with a total number of permutations $N=500$ and are bolded if $p>0.1$.
Frobenius metric AIRM metric Log-Cholesky metric
Ltoe Rtoe tongue Ltoe Rtoe tongue Ltoe Rtoe tongue
Left proposed 0.00 0.03 0.26 0.00 0.01 0.78 0.00 0.01 0.52
energy 0.01 0.05 0.48 0.01 0.48 0.51 0.02 0.30 0.54
Right proposed 0.04 0.04 0.45 0.02 0.03 0.75 0.03 0.04 0.77
energy 0.00 0.00 0.13 0.00 0.00 0.20 0.02 0.02 0.22
6.2 Age-at-death distributions
The study of human mortality as quantified by age-at-death distributions has attracted much interest over the past decades. The analysis of age-at-death distributions has provided insights into mortality patterns, lifespan variability and the factors influencing survival across different populations and time periods. The Human Mortality Database (www.mortality.org) provides annual age-at-death tables for 38 countries, with death rates recorded for each country across ages ranging from 0 to 110. By applying basic smoothing techniques to the yearly life tables available in this database, one may obtain smooth age-at-death density distributions indexed by both country and calendar year. We focus on 33 countries with available data spanning the years 1983 to 2018 and employ the conditional framework detailed in Section 4, where the data are considered to be triplets $(X,Y,Z)$. Here $X$ stands for the male age-at-death distribution for a given country at calendar year $Z$, and $Y$ the corresponding age-at-death distribution for females for the same country and year. We employ the Wasserstein-2 metric (16) in the space of one-dimensional distributions considered here, as it is particularly sensitive to horizontal shifts in distributions that are of special interest.
We apply the proposed conditional profile association $\Delta_{XY}(z)$ in (12) for this analysis, aiming to shed light on the changing nature of the relationship between male and female age-at-death distributions across different calendar years. The conditional profile association serves to quantify the degree of dependence between the two distributions as a function of calendar time $z$, where a smaller value of $\Delta_{XY}(z)$ indicate that male and female death rate distributions are relatively less dependent. dependent, while larger values indicate stronger association and a higher degree of dependence.
We divide the 33 countries into two groups: Eastern European countries, which include Bulgaria, Belarus, Czechia, Estonia, Hungary, Lithuania, Latvia, Poland, Slovakia, and Slovenia; and other countries that can be broadly characterized as Western/Asian and include Australia, Austria, Belgium, Canada, Switzerland, Denmark, Spain, Finland, France, England and Wales, Greece, Ireland, Iceland, Italy, Japan, Luxembourg, the Netherlands, Norway, New Zealand, Portugal, Sweden, Taiwan and the United States. The results are in Figure 4, which illustrates the estimated conditional profile association as a function of calendar year for all countries combined, as well as separately for Eastern European and Western/Asian countries.
The left panel, representing all countries, exhibits a generally decreasing trend over time, indicating a gradual reduction in the association between male and female age-at-death distributions over the 25 years covered in this analysis. The middle panel, which focuses on Eastern European countries only, shows a noticeable peak in the early 1990s, suggesting a period of heightened dependence between male and female mortality distributions, potentially due to major socio-political and economic changes in the region during that time that had a strong effect on both female and male age-at-death distributions. After this peak, the association declines but remains relatively stable. In contrast, the right panel, which represents Western/Asian countries, shows a consistently lower conditional profile association compared to both all countries combined and Eastern European countries. The association in Western countries also follows a steady downward trend, suggesting an increasing independence between male and female mortality distributions over time. This lower and declining association seems to indicate an increasing decoupling between male and female mortality over time that is much more expressed for Western/Asian countries but is much less evident for Eastern European countries. The clear disparity between Western and Eastern European countries highlights the varying demographic, political and health-related forces that influence the relationship between male and female mortality.
Figure 4: Estimated conditional profile associations as a function of calendar year for distributional data that correspond to male and female age-at-death distributions across 33 countries. Conditional profile associations are shown for all countries (left panel) and separately for Eastern European countries (middle panel) and Western/Asian countries (right panel). The associations are rescaled to lie between 0 and 1.
7 Discussion
In this paper, we introduce the profile association, a measure designed to quantify associations between random objects. The proposed association is based on distance profiles, which serve as bridges connecting the distributions of random objects, the classical Hoeffding’s D statistic, and the proposed association measure. Utilizing this framework, we develop an independence testing procedure for random objects, demonstrating superior performance compared to existing methods in numerical studies. We further extend the profile association and the associated testing framework to conditional settings, in which random objects are paired with Euclidean predictors. This extension provides a useful tool for exploring relationships between complex object responses and Euclidean covariates, with immediate practical relevance. Beyond independence testing, the profile association framework has potential for broader statistical applications, including causal inference, change-point detection, and feature selection, thus opening promising avenues for future research.
Causal inference has emerged as an important field in statistics, with significant applications across various scientific disciplines. One key challenge in causal inference is detecting confounding contributions, which represent spurious associations arising from the influence of a confounder on two outcomes of interest. A simple conceptual tool for understanding confounding for Euclidean data is the classical law of total covariance,
$\text{Cov}(X,Y)=\mathbb{E}[\text{Cov}(X,Y\mid Z)]+\text{Cov}(\mathbb{E}[X\mid Z],\mathbb{E}[Y\mid Z]),$
which decomposes the conditional covariance between random variables $X$ and $Y$ in the presence of a conditioning variable $Z$. Here $\text{Cov}(\mathbb{E}[X\mid Z],\mathbb{E}[Y\mid Z])$ is the contribution to the covariance explained by $Z$, and thus possibly due to a confounding effectof $Z$, whereas $\text{Cov}(\mathbb{E}[X\mid Z],\mathbb{E}[Y\mid Z])$ is the part that is not explained by $Z$ in the context of a linear framework.
An interesting question for future research is whether by utilizing the proposed conditional profile association one can extend such a separation of the effect that can be attributed to $Z$ and the remaining association that cannot be explained by $Z$ to scenarios where $X$ and $Y$ are random objects and a linear framework does not apply, due to the lack of vector operations in non-Euclidean spaces. Considering the mortality example discussed in Section 6.2, where $(X_{i}$ and $Y_{i}$ are female and male age-at-death distributions, while $Z_{i}$ is a Euclidean covariate, a preliminary attempt to shed some light on this is the following three-step procedure: first compute the unconditional profile association $\Delta(X,Y)$, using all available data pairs $\{(X_{i},Y_{i})\}_{i}$; second, replace the conditional expectations $\mathbb{E}[X\mid Z]$, $\mathbb{E}[Y\mid Z])$ in the law of total covariance by Fréchet regressions (Petersen and Müller, 2019) to obtain conditional Fréchet means $\mathbb{E}_{\oplus}[X\mid Z]$ and $\mathbb{E}_{\oplus}[Y\mid Z]$ using data $\{(X_{i},Z_{i})\}_{i}$ and $\{(Y_{i},Z_{i})\}_{i}$ respectively; third, compute $\Delta(\mathbb{E}_{\oplus}[X\mid Z],\mathbb{E}_{\oplus}[Y\mid Z])$ by plugging in the results from the Fréchet regression step.
Applying these considerations for the human mortality data for all countries combined yields $0.3630\Delta(X,Y)=$ 0.36 for the estimate of the overall profile association that ignores calendar year, but $30\Delta(\mathbb{E}[_{\oplus}X\mid Z],\mathbb{E}[_{\oplus}Y\mid Z])=0.97$ for the estimate of the part of the association explained by $Z$. Here we introduce the multiplier $30$ to normalize the association values so they lie in the interval $[0,1]$. Thus when calendar year is ignored, the observed profile association between male and female age-at-death distributions is much weaker than the value obtained when one first uses the calendar year as predictor. This discussion is not meant to be conclusive but rather to showcase the potential for future statistical research on theory and applications of conditional profile association for random objects.
Acknowledgments
This research was supported in part by NSF DMS-2310450. We thank Professor Russell Lyons, Indiana University, for important comments.
Appendix
Appendix A Metric spaces of strong negative type
In metric geometry and probability theory, the concept of negative type generalizes the idea of negative definite matrices from linear algebra and serves as a framework for analyzing how distances in a metric space interact with certain functional inequalities. Here we give a brief introduction to the definition and major properties of metric spaces of negative type. A detailed discussion of the properties of such spaces and additional examples can be found in Lyons (2013).
A metric space $(\mathcal{M},d)$ is said to be of negative type if, for any finite collection of points $\omega_{1},\ldots,\omega_{n}\in\mathcal{M}$ and real numbers $\alpha_{1},\ldots,\alpha_{n}$ such that $\sum_{i=1}^{n}\alpha_{i}=0$,
$\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{i}\alpha_{j}d(\omega_{i},\omega_{j})\leq 0.$
Examples of metric space of negative types include $\mathcal{L}_{p}$ spaces equipped with the metric $d(\omega_{1},\omega_{2})=\|\omega_{1}-\omega_{2}\|_{p}^{\min\{1,p\}}$ for $0<p\leq 2$, spheres with the geodesic distance and all real and complex hyperbolic spaces (Meckes, 2013). Importantly, $(\mathcal{M},d)$ is of negative type if and only if there is a Hilbert space $\mathcal{H}$ and an injective map ϕ::italic-$\phi:\mathcal{M}\mapsto\mathcal{H}$ such that for all $\omega,\omega^{\prime}\in\mathcal{M}$, $d(\omega,\omega^{\prime})=\|\phi(\omega)-\phi(\omega^{\prime})\|^{2}$. This equivalent definition was discovered by Schoenberg (1937, 1938), and it is the key to prove that distance profiles uniquely characterize the distribution of random objects. It also provides the theoretical foundation for numerous other methods in statistics and machine learning.
A metric space of strong negative type is a concept that additionally relates to the underlying probability measure in the metric measure space and extends the negative type property beyond the consideration of finite collections of points. Specifically, a metric space $(\mathcal{M},d)$ is said to be of strong negative type if the condition
$D(\mu,\nu):=\iint d(\omega,\omega^{\prime})\,d(\mu-\nu)^{2}(\omega,\omega^{\prime})\leq 0$
holds for all distinct Borel probability measures $\mu$ and $\nu$ on $\mathcal{M}$ with finite first moments, with equality if and only if $\mu=\nu$. The functional $D(\mu,\nu)$ is known as the energy distance between $\mu$ and $\nu$, and it generalizes the notion of a metric on the space of probability measures on the metric space $\mathcal{M}$. The strong negative type property ensures that the energy distance is zero if and only if $\mu=\nu$, meaning that the metric structure on the space allows for the identification of probability measures based solely on pairwise distances.

### v3

@@
 Notable examples for such non-traditional and non-Euclidean data are distribution-valued data and Riemannian manifold-valued data, which have numerous applications in fields such as economics (Gunsilius, 2023), biology (Bolstad et al., 2003), brain imaging (Lila et al., 2016; Lila and Aston, 2020) and geography (Mitchell et al., 2025).
-For data residing in spaces that are endowed with a (pseudo-)Riemannian geometry, a common approach is to apply transformations, such as the Riemannian log map, to project the data into a linear tangent space (Petersen and Müller, 2016; Bigot et al., 2017; Lin and Yao, 2019; Chen et al., 2023).
+For data residing in spaces that are endowed with a (pseudo-)Riemannian geometry (O’Neill, 1983), a common approach is to apply transformations, such as the Riemannian log map, to project the data into a linear tangent space (Petersen and Müller, 2016; Bigot et al., 2017; Lin and Yao, 2019; Chen et al., 2023).
 If such a linearizing map is available, one may adopt traditional methodologies for data in Euclidean spaces such as quantifications for association of multivariate or functional data with their usual Euclidean (or $\mathcal{L}^{2}$) metrics.
@@
 Such intrinsic approaches that directly confront the absence of linear operations have gained much interest in recent studies (Dubey and Müller, 2020; Zhu and Müller, 2023; Schötz, 2025; Jiang et al., 2024).
-Distance profiles (Dubey et al., 2024; Wang et al., 2024) corresponding to a class of one-dimensional distribution functions that characterize the distribution of random objects under mild conditions have proved instrumental for predictive conformal inference (Zhou and Müller, 2024a).
+Distance profiles (Dubey et al., 2024; Wang et al., 2024) corresponding to a class of one-dimensional distribution functions that characterize the distribution of random objects under mild conditions have proved instrumental for predictive conformal inference (Zhou and Müller, 2025).
 In this paper, we consider the setting where $X$ and $Y$ are random objects that may be situated in different metric spaces and develop a unified framework to quantify the association between $X$ and $Y$.
@@
 The analysis of the empirical profile association relies on a connection between the Hoeffding D statistic and the proposed profile association, which makes it possible to devise a consistent permutation test for testing the independence between $X$ and $Y$ and to derive its power properties, the profile independence test.
-Popular alternatives to obtain inference with distance-based tests include the energy test and the d-variable Hilbert-Schmidt Independence Criterion (dHSIC), which have been extended to non-Euclidean data (Székely et al., 2007; Lyons, 2013; Sejdinovic et al., 2013) and we include these methods in our numerical comparisons.
+Popular alternatives to obtain inference with distance-based tests include the energy test and ball divergence test, which have been extended to non-Euclidean data (Székely et al., 2007; Lyons, 2013; Sejdinovic et al., 2013; Wang et al., 2024; Gao et al., 2025) and we include these methods in our numerical comparisons.
+Testing independence between non-Euclidean data has become an active area of research with important applications across many interdisciplinary fields.
 For random elements taking values in Banach spaces, Pan et al.
-(2020) introduced an independence measure based on the concept of ball covariance, a concept closely related to distance profiles.
-However, the proof of the independence-zero equivalence property of ball covariance in Pan et al.
-(2020) depends on the linear structure of the Banach space.
-In practical applications, it is uncommon to encounter data that reside exclusively in a Banach space without also belonging to a Hilbert space equipped with an inner product, where one has available an array of methods that utilize the availability of linear operations (Lee et al., 2020; Lundborg et al., 2022).
-We emphasize that our proposed approach to quantify and infer dependence also applies to high-dimensional and functional data equipped with the Euclidean or $L^{2}$ metric, as it is universally applicable across all metric spaces.
-We note that recently Wang et al.
-(2024) extended the concept of ball covariance to metric spaces and this paper also features a subsection on its application for independence testing.
-However, since independence testing is only a secondary focus of this paper, the independence testing framework for random objects as presented remains incomplete, as the consistency properties along with the details and theoretical guarantees for the power of the resulting permutation implementation have not been established.
-Moreover, simulation studies demonstrate that our proposed profile independence test outperforms the ball covariance test across all settings.
+(2020) introduced an independence measure based on the concept of ball covariance, which is closely related to distance profiles.
+However, their proof of the independence-zero equivalence relies on the linear structure of Banach spaces.
+In practice, it is uncommon to encounter data that reside exclusively in Banach spaces without also belonging to a Hilbert space, where one can leverage inner product structures and apply a wide range of linear methods (Lee et al., 2020; Lundborg et al., 2022).
+Recently, Wang et al.
+(2024) extended the notion of ball covariance to general metric spaces and included a subsection on its application to independence testing, along with an analysis of the asymptotic distribution of the test statistic under the null hypothesis.
+However, the overall framework for independence testing with random objects remains incomplete: key properties such as the consistency of the test and the power analysis of the permutation-based implementation have not yet been fully established.
+We systematically study the independence testing problem for random objects based on distance profiles.
+Intrinsically, both the metric distribution functions (MDF) Wang et al.
+(2024) and the distance profiles Dubey et al.
+(2024) rely on distribution functions that characterize probabilistic information about distances in a metric space.
+In the context of independence testing, there are at least three advantages of using distance profiles over MDF.
+First, the framework of distance profiles naturally connects the independence of random objects with the independence of a class of one-dimensional random variables and satisfies a factorization property (see Proposition 1 in Section 2).
+This connection bridges the gap between dependence structures of complex objects and those of univariate random variables, allowing one to apply a wide range of existing tests for univariate independence.
+Second, because distance profiles are univariate distribution functions, it is straightforward to relate them to Hoeffding’s $D$ statistic (Hoeffding, 1948), enabling the construction of interpretable and computationally efficient test statistics.
+Third, the assumptions required for MDF and distance profiles to fully characterize probability measures on metric spaces are different.
+In Wang et al.
+(2024), it relies on the directionally $(\epsilon,\eta,L)$-limited condition (Federer, 2014), which is somewhat abstract and often difficult to verify, particularly in infinite-dimensional spaces such as Wasserstein spaces.
+In contrast, under the distance profile framework, this one-to-one correspondence holds under mild assumptions on either the metric space or the probability measure, making the approach more flexible and broadly applicable.
 Very recently and independently of our work, another study on testing mutual independence between random objects was posted on arXiv (Chen and Dubey, 2025).
-While this work also utilizes distance profiles, it does not discuss the independence-zero equivalence property, and the proposed test statistic is based on estimated distance profiles.
-Our testing framework was derived simultaneously with this work, establishes the independence-zero equivalence, provides details for the dependence-1 equivalence, is based on a different Hoeffding D-statistic-based test statistic, employs different proof techniques and utilizes a half-permutation test for the study of consistency and power.
-Furthermore, we extend our approach to the important case of testing and quantifying conditional associations, when conditioning on a Euclidean random vector.
+While this work also utilizes distance profiles, its proposed test statistic is based on estimated distance profiles.
+Our testing framework, developed simultaneously but independently, differs in several key aspects.
+It is based on a Hoeffding’s $D$-statistic-type test, establishes the dependence-1 equivalence, employs different proof techniques and employs a half permutation test to study consistency and power.
+Furthermore, we extend our approach to the important problem of testing and quantifying conditional associations, conditioning on a Euclidean random vector.
 The quantification of conditional associations and inference is motivated by the fact that many real-world datasets involve complex random objects along with Euclidean predictors, such as age, time or experimental conditions.
-Understanding conditional relationships then enables researchers to model, predict, and interpret such data more effectively, especially in fields such as medical imaging, neuroscience, genetics, and environmental science.
-Specifically, we introduce the notion of conditional profile association to quantify the dependence between random objects $X$ and $Y$ given a Euclidean predictor $Z$ We showcase the application of conditional profile association for human longevity, where we study the relationship between the age-at-death distributions of males and females across different calendar years for various countries in Section 6.2.
+Understanding conditional relationships enables researchers to model, predict, and interpret such data effectively, especially in fields such as medical imaging, neuroscience, genetics or environmental science.
+Specifically, we introduce the notion of conditional profile association to quantify the dependence between random objects $X$ and $Y$ given a Euclidean predictor $Z$.
 Using estimated joint and marginal distance profile functions, we also propose a statistic to carry out a conditional profile independence test to infer whether $X$ and $Y$ are conditionally independent and establish the uniform convergence of the estimated joint and marginal conditional distance profiles, which then implies the uniform convergence of conditional profile association.
 We also study the asymptotic distribution of the test statistic for the test for conditional independence under the null hypothesis and its implementation with a permutation test.
+We evaluate the finite-sample performance of the proposed conditional test through simulation studies and an application to human mortality data, where we examine conditional profile associations between male and female age-at-death distributions across different calendar years (see Section 6.2).
 In summary, in this paper we introduce
@@
 2 Profile association for random objects
-We state here some basic assumptions and notations.
-In the following, $X$ and $Y$ are random elements that take values in totally bounded, separable metric spaces $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with associated metrics $d_{1}(\cdot,\cdot)$ and $d_{2}(\cdot,\cdot)$, respectively.
+We state here basic assumptions and notations.
+In the following, $X$ and $Y$ are random elements that take values in totally bounded separable metric spaces $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with metrics $d_{1}(\cdot,\cdot)$ and $d_{2}(\cdot,\cdot)$, respectively.
 We refer to such random elements as random objects.
 We denote by $(S,\mathcal{S},\mathbb{P})$ an underlying probability space, where $\mathcal{S}$ is the Borel $\sigma$-algebra on $S$ and $\mathbb{P}$ is a probability measure, and by $\mu_{XY}$ the joint measure of $X$ and $Y$, where $\mu_{XY}(A\times B)=\mathbb{P}(s\in S:X(s)\in A\text{ and }Y(s)\in B)$ for any Borel measurable sets $A\subset\mathcal{M}_{1}$ and $B\subset\mathcal{M}_{2}$.
-We further assume that the marginal distributions of $X$ and $Y$, denoted by $\mu_{X}$ and $\mu_{Y}$, have finite first-order moments, i.e., $\int d_{1}(\omega_{1},x)\,\mathrm{d}\mu_{X}(x)<\infty$ and $\int d_{2}(\omega_{2},x)\,\mathrm{d}\mu_{Y}(x)<\infty$ for some (and thus any) $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
-We are interested in quantifying the association between $X$ and $Y$, and especially in determining whether $X$ and $Y$ are independent, which is equivalent to $\mu_{XY}=\mu_{X}\times\mu_{Y}$.
-The following proposition states a necessary and sufficient condition for $X$ and $Y$ to be independent.
+The marginal distributions of $X$ and $Y$, denoted by $\mu_{X}$ and $\mu_{Y}$, are assumed to have finite first-order moments, i.e., $\int d_{1}(\omega_{1},x)\,\mathrm{d}\mu_{X}(x)<\infty$ and $\int d_{2}(\omega_{2},x)\,\mathrm{d}\mu_{Y}(x)<\infty$ for some (and thus any) $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
+We aim to quantify the association between $X$ and $Y$ and to determine whether $X$ and $Y$ are independent, which is equivalent to $\mu_{XY}=\mu_{X}\times\mu_{Y}$.
 In general metric spaces, linear operations like addition, subtraction and inner product that are fundamental to conventional statistical inference for Euclidean data are not available.
 Furthermore, in some metric spaces like $\mathcal{L}^{2}{[0,1]}$, the absence of a well-defined Lebesgue measure renders density-based inference methods infeasible.
-Since the only structure available in general metric spaces is the distance between two elements, developing intrinsic statistical methodologies that align with the underlying geometry requires leveraging these distances.
-The following proposition provides a sufficient condition under which the independence of probability measures is equivalent to the independence of distances between random elements and every point in the metric space.
+Since the only structure available in general metric spaces is the distance between two elements, developing intrinsic statistical methodologies that align with the underlying geometry requires working directly with distances.
+We further note that to develop statistical methods for general metric spaces, one needs to restrict attention to a subclass of probability measures that are fully determined by their values on closed balls.
+Specifically, this means that the values $\mu(\overline{\mathcal{B}}(\omega,r))$ for all points $\omega\in\mathcal{M}$ and all radii $r>0$ uniquely determines the measure $\mu$ on all Borel sets, where $\overline{\mathcal{B}}(\omega,r):=\left\{\omega^{\prime}\in\mathcal{M}:d(\omega,\omega^{\prime})\leq r\right\}$ denotes the closed ball centered at $\omega$ with radius $r$.
+This property is commonly referred to as the one-to-one correspondence and was first studied by Christensen (1970), Davies (1971), and Hoffmann-Jørgensen (1975).
+More recently, it has been investigated by Blanche et al.
+(2016), with related applications in statistics and probability discussed in Lyons (2013), Dubey et al.
+(2024), and Chen and Dubey (2025).
+Distance profiles (Dubey et al., 2024) provide a natural framework for studying the one-to-one correspondence and for developing inference tools for random objects by establishing an intrinsic connection between such objects and a class of one-dimensional distribution functions.
+This general framework has recently been applied for the construction of two-sample tests (Dubey et al., 2024; Wang et al., 2024) and for conformal prediction (Zhou and Müller, 2025).
+Specifically, for each $\omega_{1}\in\mathcal{M}_{1}$, the marginal distance profile with respect to $\mu_{X}$ is defined as
+$F^{X}_{\omega_{1}}(u)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u\right),\,\,\,u\geq 0,$ (1)
+and the marginal distance profiles $F^{Y}_{\omega_{2}}(u)$ for each $\omega_{2}\in\mathcal{M}_{2}$ with respect to $\mu_{Y}$ are defined analogously.
+To quantify the dependence and characterize independence between $X$ and $Y$, we extend the marginal distance profiles in (1) to a two-dimensional version by introducing joint distance profiles for each pair $(\omega_{1},\omega_{2})\in\mathcal{M}_{1}\times\mathcal{M}_{2}$ with respect to $\mu_{XY}$:
+$F_{\omega_{1},\omega_{2}}^{XY}(u,v)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u,\,d_{2}(Y,\omega_{2})\leq v\right),\quad\,u,v\geq 0.$ (2)
+The concept of joint distance profiles is a natural extension of distance profiles and has also been proposed independently in Chen and Dubey (2025).
+The following factorization proposition is a fundamental result that enables us to transform the independence testing problem for random objects into testing a class of two-dimensional vectors.
 Proposition 1.
-If the probability measures $\mu_{X}$, $\mu_{Y}$, and $\mu_{XY}$ are fully determined by their values on closed balls, then $\mu_{XY}=\mu_{X}\times\mu_{Y}$ if and only if $d_{1}(X,\omega_{1})$ is independent of $d_{2}(Y,\omega_{2})$ for all $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
-A probability measure $\mu$ is said to be “fully determined by its values on closed balls” if knowing the values $\mu(\overline{\mathcal{B}}(\omega,r))$ for all points $\omega\in\mathcal{M}$ and all radii $r>0$ is sufficient to uniquely specify $\mu$ on all Borel sets.
-This problem was first proposed and studied by Christensen (1970), Davies (1971), and Hoffmann-Jørgensen (1975), and has been recently investigated by Blanche et al.
-(2016), along with applications in statistics and probability discussed in Lyons (2013), Dubey et al.
-(2024), and Chen and Dubey (2025).
-The following proposition provides two sufficient conditions under which the condition in Proposition 1 holds.
-A probability measure $\mu$ on a metric space $(\mathcal{M},d)$ is said to be doubling if there exists a constant $L_{\mu}\geq 0$ such that
-$\mu\left\{\overline{\mathcal{B}}(\omega,2r)\right\}\leq L_{\mu}\mu\left\{\overline{\mathcal{B}}(\omega,r)\right\}\text{ for all }r>0\text{ and }\omega\in\mathrm{supp}(\mu),$ (1)
-where ¯$\overline{\mathcal{B}}(\omega,r):=\left\{\omega^{\prime}\in\mathcal{M}:d(\omega,\omega^{\prime})\leq r\right\}$ is the closed ball centered at $\omega$ with radius $r$, and $\mathrm{supp}(\mu):=\left\{\omega^{\prime}\in\mathcal{M}:\mu(\{\omega^{\prime}\})>0\right\}$ is the support of $\mu$.
-The concept of a doubling measure in metric spaces was first introduced by Beurling and Ahlfors (1956) and has since become a fundamental notion in analysis on metric spaces, including various areas such as harmonic analysis and geometric measure theory (Heinonen, 2001).
-Proposition 2.
-Proposition 1 holds provided that either of the following assumptions is satisfied:
+Suppose that either of the following assumptions holds:
   • (M0).
@@
   • (P0).
-    The measures $\mu_{X}$, $\mu_{Y}$ and $\mu_{XY}$ satisfy the doubling condition (1).
-Assumption M0 pertains to the structural properties of the underlying metric spaces.
-The proof is based on Proposition 3.5, Proposition 3.7, and Lemma 3.8 in Lyons (2013).
-Preliminary background on metric spaces of strong negative type is provided in Appendix A, with additional details and illustrative examples available in Klebanov et al.
+    The measures $\mu_{X}$, $\mu_{Y}$, and $\mu_{XY}$ are assumed to satisfy the doubling condition:
+    $\mu\left\{\overline{\mathcal{B}}(\omega,2r)\right\}\leq L_{\mu}\mu\left\{\overline{\mathcal{B}}(\omega,r)\right\}\text{ for all }r>0\text{ and }\omega\in\mathrm{supp}(\mu),$ (3)
+    for $\mu=\mu_{X},\mu_{Y},$ and $\mu_{XY}$, where $\mathrm{supp}(\mu):=\left\{\omega^{\prime}\in\mathcal{M}:\mu(\{\omega^{\prime}\})>0\right\}$ denotes the support of $\mu$ and $L_{\mu}$ is a positive constant.
+Then the random objects $X$ and $Y$ are independent if and only if
+$F_{\omega_{1},\omega_{2}}^{XY}(t,s)=F_{\omega_{1}}^{X}(t)\cdot F_{\omega_{2}}^{Y}(s),\quad\text{for all }t,s\geq 0,\ \omega_{1}\in\mathcal{M}_{1},\ \omega_{2}\in\mathcal{M}_{2}.$
+Assumption M0 relates to the structural properties of the underlying metric spaces.
+Under this assumption, Corollary 3.20 of Lyons (2013) shows that the product space $\mathcal{M}_{1}\times\mathcal{M}_{2}$ is also of strong negative type.
+A brief review of the strong negative type assumption is provided in Supplement, with additional details and examples available in Klebanov et al.
 (2005) and Lyons (2013).
-Assumption P0 concerns the properties of the associated probability measures and was employed in recent work Chen and Dubey (2025) to study mutual independence in metric spaces.
-Proposition 1 and Proposition 2 establish a connection between the independence of probability measures and the independence of distances of random elements, providing the basis for statistical inference for association and independence in metric spaces.
-Our approach utilizes distance profiles (Dubey et al., 2024), where for each $\omega_{1}\in\mathcal{M}_{1}$, the marginal distance profile with respect to $\mu_{X}$ is defined as
-$F^{X}_{\omega_{1}}(u)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u\right),\,\,\,u\geq 0,\,\,\omega_{1}\in\mathcal{M}_{1},$ (2)
-and marginal distance profiles $F^{Y}_{\omega_{2}}(u)$ for each $\omega_{2}\in\mathcal{M}_{2}$ with respect to $\mu_{Y}$ are defined analogously.
-Distance profiles establish an intrinsic link between random objects and the one-dimensional distribution functions $F_{\omega}$ and provide a feasible framework for developing inference tools for random objects.
-This framework has been previously applied for the construction of two-sample tests (Dubey et al., 2024; Wang et al., 2024) and conformal prediction (Zhou and Müller, 2024a).
-To measure the independence between $X$ and $Y$, we adopt the idea to extend (2) to the two-dimensional case by considering joint distance profiles for pairs of elements $(\omega_{1},\omega_{2})$,   $\omega_{1}\in\mathcal{M}_{1},\omega_{2}\in\mathcal{M}_{2}$, with respect to $\mu_{XY}$,
-$F_{\omega_{1},\omega_{2}}^{XY}(u,v)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u,\,d_{1}(Y,\omega_{2})\leq v\right),\quad\,u,v\geq 0,\,\,\,\omega_{1}\in\mathcal{M}_{1},\,\,\omega_{2}\in\mathcal{M}_{2}.$ (3)
-Proposition 2 then implies the following result, which relates the independence between random objects to their corresponding distance profiles.
-Proposition 3.
-Under either (M0) or (P0), the random objects $X$ and $Y$ are independent if and only if
-$F_{\omega_{1},\omega_{2}}^{XY}(t,s)=F_{\omega_{1}}^{X}(t)F_{\omega_{2}}^{Y}(s),$
-for all $t,s\geq 0$, $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
-Proposition 3 enables us to transform the independence testing problem for random objects into testing a class of two-dimensional vectors $\{(d_{1}(X,\omega_{1}),d_{2}(Y,\omega_{2}))^{\top}\}$ indexed by $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
-Extensive research has been conducted on testing the independence of vector-valued random variables $X$ and $Y$ by quantifying the discrepancy between their joint measure and the product of their marginal measures.
+Assumption P0 imposes conditions on the probability measures and was adopted in recent work by Chen and Dubey (2025) to study mutual independence in metric spaces.
+The concept of a doubling measure in metric spaces was first introduced by Beurling and Ahlfors (1956) and has since become a fundamental notion for metric measure spaces, harmonic analysis and geometric measure theory (Heinonen, 2001).
+The significance of Proposition 1 is that it reduces the complex independence testing problem for random objects to the simpler problem of studying the dependence of two-dimensional vectors $\{(d_{1}(X,\omega_{1}),d_{2}(Y,\omega_{2}))^{\top}\}$, indexed by $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$, where one can draw on extensive existing research.
+A standard approach for testing the independence of vector-valued random variables $X$ and $Y$ is to quantify the discrepancy between their joint measure and the product of their marginal measures.
 For scalar random variables $X$ and $Y$ with joint distribution function $F(x,y)$, in a seminal paper Hoeffding (1948) introduced the functional
 $\Delta(F)=\iint\{F(x,y)-F(x,\infty)F(\infty,y)\}^{2}\,\mathrm{d}F(x,y),$
-which corresponds to a type of profile association between $X$ and $Y$. When $F(x,y)$ has a continuous density function, $X$ and $Y$ are independent if and only if $\Delta(F)=0$.
+which quantifies the association between $X$ and $Y$. When $F(x,y)$ has a continuous density function, $X$ and $Y$ are independent if and only if $\Delta(F)=0$.
 When $X$ and $Y$ are random objects, one may analogously define their profile association as
 $\Delta(X,Y):=\mathbb{E}_{X^{\prime},Y^{\prime}}\left[\iint\{F_{X^{\prime},Y^{\prime}}^{XY}(u,v)-F_{X^{\prime}}^{X}(u)F_{Y^{\prime}}^{Y}(v)\}^{2}{\rm d}F^{XY}_{X^{\prime},Y^{\prime}}(u,v)\right],$ (4)
-where $(X^{\prime},Y^{\prime})$ is an independent copy of $(X,Y)$, and the random distance profiles $F_{X^{\prime},Y^{\prime}}^{XY}(u,v)$, $F_{X^{\prime}}^{X}(u)$, and $F_{Y^{\prime}}^{Y}(v)$ are defined as:
-$F_{X^{\prime},Y^{\prime}}^{XY}(u,v):=\mathbb{P}_{X,Y}\left(d_{1}(X,X^{\prime})\leq u,\,d_{2}(Y,Y^{\prime})\leq v\right),$
-$F_{X^{\prime}}^{X}(u):=\mathbb{P}_{X}\left(d_{1}(X,X^{\prime})\leq u\right),\quad F_{Y^{\prime}}^{Y}(v):=\mathbb{P}_{Y}\left(d_{2}(Y,Y^{\prime})\leq v\right),$
+where $(X^{\prime},Y^{\prime})$ is an independent copy of $(X,Y)$, and the random distance profiles $F_{X^{\prime},Y^{\prime}}^{XY}(u,v)$, $F_{X^{\prime}}^{X}(u)$, and $F_{Y^{\prime}}^{Y}(v)$ are defined as
+$F_{X^{\prime},Y^{\prime}}^{XY}(u,v)=\mathbb{P}_{X,Y}\left(d_{1}(X,X^{\prime})\leq u,\,d_{2}(Y,Y^{\prime})\leq v\right),$
+$F_{X^{\prime}}^{X}(u)=\mathbb{P}_{X}\left(d_{1}(X,X^{\prime})\leq u\right),\quad F_{Y^{\prime}}^{Y}(v)=\mathbb{P}_{Y}\left(d_{2}(Y,Y^{\prime})\leq v\right),$
 for all $u,v\geq 0$.
 Throughout this paper, the subscript in $\mathbb{E}_{X}$ (or $\mathbb{P}_{X}$) indicates that the expectation (or probability) is taken over the random variable(s) specified in the subscript only.
-For a random object $X$, i.e., a metric space-valued random variable, we denote by $X(s)\in\mathcal{M}$ the value it assumes in the metric space $\mathcal{M}$ for a specific argument $s$ that is an element of the underlying probability space $S$.
+For a random object $X$, i.e., a metric space-valued random variable, we denote by $X(s)\in\mathcal{M}_{1}$ the value it assumes in the metric space $\mathcal{M}_{1}$ for a specific argument $s$, which is an element of the underlying probability space $S$.
 Theorem 1.
@@
 Specifically, when $\mathcal{M}_{1}=\mathcal{M}_{2}=\mathbb{R}$, this assumption can be derived from the condition in Hoeffding (1948), which requires that the pair $(X,Y)$ has a continuous joint density function.
-The first statement of Theorem 1 shows that the proposed profile association measure satisfies both I-consistency and D-consistency; see Weihs et al.
-(2018):
-  • I-consistency. If $X$ and $Y$ are independent, the profile association measure is zero.
-  • D-consistency. If $X$ and $Y$ are dependent, the profile association measure is nonzero.
-A similar independence-zero equivalence property has been established for ball covariance (Pan et al., 2020), which was proposed to test independence between Banach space-valued random elements (while we consider here random objects which take values in more general metric spaces).
-The proof of Theorem 1 relies on a mapping between negative-type metric spaces and Hilbert spaces (Schoenberg, 1937, 1938; Lyons, 2013).
+The first statement of Theorem 1 shows that the proposed profile association measure satisfies both I-consistency and D-consistency (Weihs et al., 2018):
+  • I-consistency.
+    If $X$ and $Y$ are independent, the profile association measure is zero.
+  • D-consistency.
+    If $X$ and $Y$ are dependent, the profile association measure is nonzero.
+A similar independence-zero equivalence property has been established for ball covariance (Pan et al., 2020), which was proposed to test independence for the special case of Banach (and thus vector-)space-valued random elements.
+The proof of Theorem 1 utlizes a well-known mapping between negative-type metric spaces and Hilbert spaces (Schoenberg, 1937, 1938; Lyons, 2013).
 Theorem 1 states that $\Delta(X,Y)$ achieves its maximum value if and only if $X=Y$ almost surely when $\mathcal{M}_{1}=\mathcal{M}_{2}$.
@@
 The proposed profile association measure (4) specifically targets the quantification of association for random objects in general metric spaces.
-We note that monotone functions are not well-defined in metric spaces that lack a natural order and furthermore that alternative measures of independence often assign the same maximal association level to any monotone (or measurable) function of $X$, offering no additional insight into the degree of dependence between $Y$ and $X$.
+We note that monotone functions are not well-defined in metric spaces that lack a natural order and furthermore that alternative measures of independence often assign the same maximal association level to any monotone (or measurable) function of $X$, providing no additional insight into the degree of dependence between $Y$ and $X$.
 This limitation prevents these measures from distinguishing varying levels of dependence.
@@
 3.1 Testing independence for random objects
-We are interested in testing the hypothesis
-[…]
+We are interested in testing
+$\mathcal{H}_{0}:X\perp\!\!\!\perp Y\quad\text{versus}\quad\mathcal{H}_{1}:X\not\!\perp\!\!\!\perp Y$
 using random samples $\mathcal{D}_{\mathsf{XY}}:=\{(X_{i},Y_{i}):1\leq i\leq n\}$ drawn from the joint distribution of the random objects $(X,Y)$.
-The profile association $\Delta(X,Y)$ in (4) quantifies the strength of the relationship between random objects $X$ and $Y$ and provides the foundation for a test statistic to assess their independence.
-Consider the functions $\psi_{1}:\mathcal{M}_{1}^{\otimes 4}\mapsto\{-1,0,1\}$ defined as
+The profile association $\Delta(X,Y)$ in (4) quantifies the strength of the relationship between random objects $X$ and $Y$ and provides the foundation for the development of a test for independence.
+As discussed in Section 2, for scalar random variables $X$ and $Y$ with joint CDF $F$, Hoeffding (1948) introduced the functional $\Delta(F)=\iint\{F(x,y)-F(x,\infty)\,F(\infty,y)\}^{2}\,{\rm d}F(x,y)$ as a measure of dependence.
+This functional can be expressed in terms of indicator functions.
+Specifically, define $\psi(x_{1},x_{2},x_{3})=\mathds{1}_{\{x_{1}-x_{2}\geq 0\}}-\mathds{1}_{\{x_{1}-x_{3}\geq 0\}},$ where the indicator function $\mathds{1}_{A}=1$ if the event $A$ is true and $\mathds{1}_{A}=0$ otherwise.
+Setting $\phi(x_{1},y_{1};\ldots,x_{5},y_{5})=0.25\times\psi(x_{1},x_{2},x_{3})\psi(x_{1},x_{4},x_{5})\psi(y_{1},y_{2},y_{3})\psi(y_{1},y_{4},y_{5}),$ one can write
+$\Delta(F)=\int\cdots\int\phi(x_{1},y_{1};\ldots,x_{5},y_{5})\,{\rm d}F(x_{1},y_{1})\cdots{\rm d}F(x_{5},y_{5}).$
+When $X$ and $Y$ are random objects, we define the scalar random variables $U_{\omega_{1}}=d_{1}(X,\omega_{1})$ and $V_{\omega_{2}}=d_{2}(Y,\omega_{2})$, and denote their joint CDF by $\mathfrak{F}_{\omega_{1},\omega_{2}}$ for $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
+Using these notations, the profile association defined in (4) can be written as $\Delta(X,Y)=\mathbb{E}_{X^{\prime},Y^{\prime}}\left[\Delta\left(\mathfrak{F}_{X^{\prime},Y^{\prime}}\right)\right].$
+Inspired by this, consider the functions $\psi_{1}:\mathcal{M}_{1}^{\otimes 4}\mapsto\{-1,0,1\}$ defined as
 $\psi_{1}(x_{1},x_{2},x_{3},x_{4})=\mathds{1}_{\{d_{1}(x_{1},x_{3})\leq d_{1}(x_{1},x_{2})\}}-\mathds{1}_{\{d_{1}(x_{1},x_{4})\leq d_{1}(x_{1},x_{2})\}},\quad x_{i}\in\mathcal{M}_{1},$
@@
 $h\big((x_{1},y_{1}),\ldots,(x_{6},y_{6})\big)=$ $\frac{1}{4}\psi(x_{1},x_{2},x_{3},x_{4})\psi(x_{1},x_{2},x_{5},x_{6})$
-×$\quad\times\psi(y_{1},y_{2},y_{3},y_{4})\psi(y_{1},y_{2},y_{5},y_{6}).$
-In what follows, we use the notation $\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{6},Y_{6})\}$ to denote six independent copies of $(X,Y)$, where one has
+$\quad\times\psi(y_{1},y_{2},y_{3},y_{4})\psi(y_{1},y_{2},y_{5},y_{6}).$
+In what follows, we use the notation $\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{6},Y_{6})\}$ to denote six independent copies of $(X,Y)$. Following the previous discussion, we have
 $\mathbb{E}\Bigl[h\bigl((X_{1},Y_{1}),\ldots,(X_{6},Y_{6})\bigr)\Bigr]\;=\;\Delta(X,Y).$
-Given random samples $\mathcal{D}_{\mathsf{XY}}$, we employ a U-statistic as an estimator of $\Delta(X,Y)$, defined as
-$D_{n}(\mathcal{D}_{\mathsf{XY}})=\frac{(n-6)!}{n!}\sum_{\mathcal{P}_{6}^{n}(\bm{\alpha})}h\left((X_{\alpha_{1}},Y_{\alpha_{1}}),\ldots,(X_{\alpha_{6}},Y_{\alpha_{6}})\right),$ (5)
-where $\sum_{\mathcal{P}_{6}^{n}(\bm{\alpha})}$ demotes summation over all $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{6})^{\top}$ such that
-$\alpha_{i}=1,\ldots,n;\,\,\alpha_{i}\neq\alpha_{j}\text{ if }i\neq j,\,\,\text{ for all }\,i,j=1,2,\ldots,6.$
-We require some concepts related to the analysis of U-statistics, which were first introduced by Hoeffding (1948), with further details provided in Serfling (2009).
 Since the function $h$ is not symmetric, we first symmetrize it as follows,
 $\mathfrak{h}((x_{1},y_{1}),\ldots,(x_{6},y_{6}))=\frac{1}{6!}\sum_{\mathcal{P}_{6}^{6}(\bm{\alpha})}h((x_{\alpha_{1}},y_{\alpha_{1}}),\ldots,(x_{\alpha_{6}},y_{\alpha_{6}})),$
-where $\mathcal{P}_{6}^{6}(\bm{\alpha})$ denotes the set of all permutations of indices $\{1,2,\ldots,6\}$. For $r=1,2,\ldots,6$, let
+where $\mathcal{P}_{6}^{6}(\bm{\alpha})$ denotes the set of all permutations $\{\alpha_{1},\ldots,\alpha_{6}\}$ of $\{1,2,\ldots,6\}$.
+Given random samples $\mathcal{D}_{\mathsf{XY}}$, we employ a U-statistic as an estimator of $\Delta(X,Y)$, defined as
+$D_{n}(\mathcal{D}_{\mathsf{XY}})=\binom{n}{6}^{-1}\sum_{\mathcal{C}_{6}^{n}(\bm{\alpha})}\mathfrak{h}\left((X_{\alpha_{1}},Y_{\alpha_{1}}),\ldots,(X_{\alpha_{6}},Y_{\alpha_{6}})\right),$ (5)
+where $\sum_{\mathcal{C}_{6}^{n}(\bm{\alpha})}$ demotes summation over all $\binom{n}{6}$ combinations of $6$ distinct elements $\{\alpha_{1},\ldots,\alpha_{6}\}$ from $\{1,2,\ldots,n\}$.
+We require some concepts related to the analysis of U-statistics, which were first introduced by Hoeffding (1948), with further details provided in Serfling (2009). For $r=1,2,\ldots,6$, let
 $\mathfrak{h}_{r}((x_{1},y_{1}),\ldots,(x_{r},y_{r}))$
@@
 Assuming that $\mathbb{E}[\mathfrak{g}^{2}_{1}(X_{1},X_{2})\mathfrak{g}^{2}_{2}(Y_{1},Y_{2})]\neq 0$, we have
-[…]
+${n}(D_{n}(\mathcal{D}_{\mathsf{XY}})-\Delta(X,Y))\xrightarrow{d}{15}\sum_{j=1}^{\infty}\lambda_{j}(Z_{j}-1),$
 where $Z_{j}$ are independent $\chi_{1}^{2}$ random variables, and $\lambda_{j}$ are the eigenvalues of the linear operator $\mathscr{H}:\mathcal{L}_{\mu_{XY}}^{2}(\mathcal{M}_{1}\otimes\mathcal{M}_{2})\to\mathcal{L}_{\mu_{XY}}^{2}(\mathcal{M}_{1}\otimes\mathcal{M}_{2})$, defined by
 $\mathscr{H}(f)$ $=\iint\tilde{\mathfrak{h}}_{2}((x,y),(u,v))f(x,y){\rm d}\mu_{XY}(x,y),$
-for all $(u,v)\in\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ and $f\in\mathcal{L}_{\mu_{XY}}^{2}(\mathcal{M}_{1}\otimes\mathcal{M}_{2})$.
+for all $(u,v)\in\mathcal{M}_{1}\times\mathcal{M}_{2}$ and $f\in\mathcal{L}_{\mu_{XY}}^{2}(\mathcal{M}_{1}\times\mathcal{M}_{2})$.
 3.2 Permutation test and asymptotic results
-Since the asymptotic distribution in Theorem 2 is unknown, we use a permutation test to determine the critical values of the test statistic $D_{n}(\mathcal{D}_{\mathsf{XY}})$ when implementing the test.
+Since the asymptotic distribution in Theorem 2 is a second-order Wiener chaos with unknown parameters, we use a permutation test to determine the critical values of the test statistic $D_{n}(\mathcal{D}_{\mathsf{XY}})$ when implementing the test.
 Permuting the data and recalculating the test statistic for each permutation is a standard approach to generate a reference distribution.
-In testing independence, the typical approach is to compute replicates of the test statistic by randomly permuting the $Y$ sample while keeping the order of the $X$ sample fixed (Székely et al., 2007).
+In this subsection, we implement the proposed independence test using a permutation-based method and study its asymptotic properties, including consistency and power analysis.
+Recently, alternative approaches such as the random-lifter method (Gao et al., 2025) and spectrum-based methods (Wang et al., 2024) have been studied to handle independence testing in the presence of unknown limiting distributions.
+We include a discussion of the spectrum-based approach in the supplementary material.
+For testing independence, a common approach is to compute replicates of the test statistic by randomly permuting the $Y$ sample while keeping the order of the $X$ sample fixed (Székely et al., 2007).
 However, randomly permuting the entire $Y$ sample introduces dependence among the data pairs under the alternative $\mathcal{H}_{1}$ when $X$ and $Y$ are dependent, making it challenging to establish consistency of the critical value under $\mathcal{H}_{1}$.
-Here we adopt a half-permutation procedure introduced in Fan et al.
-(2024) that addresses this issue and does not require the splitting technique with its possible loss of power.
-Adopting this approach for our setting requires a different proof technique to obtain consistency and power bounds for the proposed profile independence test, as previous techniques are not applicable.
-We use $\mathcal{P}_{\lfloor n/2\rfloor}^{n}$ to denote the set of permutations of $\{1,2,\ldots,n\}$ with cardinality $\lfloor n/2\rfloor$.
+To address this issue, we adopt a half-permutation procedure introduced in Fan et al.
+(2024) that does not require a splitting technique with its possible loss of power.
+Adopting this approach for our setting requires a different proof technique to obtain consistency and power bounds for the proposed profile independence test, as the previous techniques are not applicable.
+We use $\mathcal{P}_{\lfloor n/2\rfloor}^{n}$ to denote the set of all subsets of $\{1,2,\ldots,n\}$ of size $\lfloor n/2\rfloor$, where each subset consists of unrepeated random samples from $\{1,\ldots,n\}$.
 Consider a sequence of i.i.d.
@@
 The permuted data corresponding to $(\Pi_{j},\sigma_{jk})$ are
-$\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}):=\{(X_{\imath_{1}},Y_{\jmath_{1}}),(X_{\imath_{2}},Y_{\jmath_{2}}),\ldots,(X_{\imath_{\lfloor n/2\rfloor}},Y_{\jmath_{\lfloor n/2\rfloor}})\}.$
-The test statistic based on permuted data $\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})$ is
-$D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})):=\frac{(\lfloor n/2\rfloor-6)!}{\lfloor n/2\rfloor!}\sum_{\mathcal{P}_{6}^{\lfloor n/2\rfloor}(\bm{\alpha})}h((X_{\imath_{\alpha_{1}}},Y_{\jmath_{\alpha_{1}}}),\ldots,(X_{\imath_{\alpha_{6}}},Y_{\jmath_{\alpha_{6}}})).$
+$\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}):=\{(X_{\imath_{1}},Y_{\jmath_{1}}),(X_{\imath_{2}},Y_{\jmath_{2}}),\ldots,(X_{\imath_{\lfloor n/2\rfloor}},Y_{\jmath_{\lfloor n/2\rfloor}})\},$
+and the test statistic based on permuted data $\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})$ is
+$D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})):=\binom{\lfloor n/2\rfloor}{6}^{-1}\sum_{\mathcal{C}_{6}^{\lfloor n/2\rfloor}(\bm{\alpha})}h((X_{\imath_{\alpha_{1}}},Y_{\jmath_{\alpha_{1}}}),\ldots,(X_{\imath_{\alpha_{6}}},Y_{\jmath_{\alpha_{6}}})).$
 Let $N=\sum_{j=1}^{B}N_{j}$ be the total number of permutation pairs $(\Pi_{j},\sigma_{jk})$.
 Since the pairs $(\Pi_{j},\sigma_{jk})$ are random elements, we can define a randomized test at level $\alpha$ by
-$\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}}):=\mathbb{P}\left(1+\sum_{(\Pi_{j},\sigma_{jk})}\mathds{1}_{\{n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})\leq\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}}\leq(1+N)\alpha\right),$ (8)
-where the probability in (8) is taken over $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and the only randomness in $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ comes from the random permutations $(\Pi_{j},\sigma_{jk})$.
+$\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}}):=\mathds{1}\left(1+\sum_{(\Pi_{j},\sigma_{jk})}\mathds{1}_{\{n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})\leq\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}}\leq(1+N)\alpha\right),$ (8)
+where the only randomness in $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ comes from the random permutations $(\Pi_{j},\sigma_{jk})$ and $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})=1$ indicates that the test rejects $H_{0}$ at level $\alpha$.
 Under the null hypothesis, $X$ and $Y$ are independent, and the elements in the data sets $\mathcal{D}_{\mathsf{XY}}$ and $\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk})$ are exchangeable.
-Consequently, $\{\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}_{j,k}$ and $n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})$ have the same asymptotic distribution by Theorem 2.
+Consequently, $\{\lfloor n/2\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\}_{j,k}$ and $n\mathcal{D}_{n}(\mathcal{D}_{\mathsf{XY}})$ have the same asymptotic distribution given by Theorem 2.
 Therefore, the probability on the right-hand side of equation (8) equals $\alpha$ for all distributions of $(X,Y)$ under the null hypothesis, which implies that $\psi_{\alpha}(\mathcal{D}_{\mathsf{XY}})$ is a valid size-$\alpha$ test.
@@
 $\hat{H}_{N}(z)=\frac{1}{N}\sum_{(\Pi_{j},\sigma_{jk})}\mathds{1}_{\left\{\lfloor\frac{n}{2}\rfloor D_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XY}}(\Pi_{j},\sigma_{jk}))\leq z\right\}}$ (9)
-for $H(z)$ based on the permuted data, where $N=\sum_{j=1}^{B}N_{j}$ is the total number of permutation pairs $(\Pi_{j},\sigma_{jk})$.
+for $H(z)$ based on the permuted data.
 Theorem 3 below establishes the consistency of this estimator, which ensures that the size of the proposed profile independence test is correctly controlled under the alternative.
@@
 Conditional independence testing is a statistical tool used to determine whether two variables are independent when conditioning on a third variable.
-It is used in a wide range of statistical applications where one studies complex relationships among variables, including causal inference (Zhang et al., 2011; Zhang et al., 2017; Peters et al., 2017), graphical models (Koller and Friedman, 2009; Williams and Mulder, 2020) and sufficient dimension reduction (Cook, 2004; Chen et al., 2015; Li, 2018).
+It is used in a wide range of statistical applications where one studies complex relationships among variables, including causal inference (Zhang et al., 2011, 2017; Peters et al., 2017), graphical models (Koller and Friedman, 2009; Williams and Mulder, 2020) and sufficient dimension reduction (Cook, 2004; Chen et al., 2015; Li, 2018).
 Many nonparametric conditional independence tests have been studied based on conditional density functions (Huang, 2010; Su and White, 2008; Banerjee et al., 2024), distribution functions (Linton and Gozalo, 1996; Zhou et al., 2020) and characteristic functions (Su and White, 2007; Wang et al., 2015; Wang and Hong, 2018).
@@
 To our knowledge, conditional independence tests for random elements in general metric spaces that do not support linear operations have remained unexplored.
-In this section, we consider a setting where $X\in\mathcal{M}_{1}$ and $Y\in\mathcal{M}_{2}$ are random objects, while the conditional variable $Z$ is Euclidean, specifically $Z\in\mathcal{Z}$, where $\mathcal{Z}$ is a compact subset of ${\mathbb{R}}^{p}$.
+We consider a setting where $X\in\mathcal{M}_{1}$ and $Y\in\mathcal{M}_{2}$ are random objects, while the conditional variable $Z$ is Euclidean, specifically $Z\in\mathcal{Z}$, where $\mathcal{Z}$ is a compact subset of ${\mathbb{R}}^{p}$.
 Our goal is to test the following hypothesis:
-[…]
+$\mathcal{H}_{0,c}:X\mid Z\perp\!\!\!\perp Y\mid Z\quad\text{versus}\quad\mathcal{H}_{1,c}:X\mid Z\not\!\perp\!\!\!\perp Y\mid Z,$ (11)
 using i.i.d.
@@
 where $F_{X^{\prime},Y^{\prime},Z^{\prime}}^{XY}$, $F_{X^{\prime},Z^{\prime}}^{X}$ and $F_{Y^{\prime},Z^{\prime}}^{Y}$ are given by
-$F_{X^{\prime},Y^{\prime},Z^{\prime}}^{XY}(u,v):=\mathbb{P}_{X,Y,Z}\left(d(X,X^{\prime})\leq u,\,d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right),$
-$F_{X^{\prime},Z^{\prime}}^{X}(u):=\mathbb{P}_{X,Z}\left(d(X,X^{\prime})\leq u\mid Z=Z^{\prime}\right),\,F_{Y^{\prime},Z^{\prime}}^{Y}(v):=\mathbb{P}_{Y,Z}\left(d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right).$
+$F_{X^{\prime},Y^{\prime},Z^{\prime}}^{XY}(u,v)=\mathbb{P}_{X,Y,Z}\left(d(X,X^{\prime})\leq u,\,d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right),$
+$F_{X^{\prime},Z^{\prime}}^{X}(u)=\mathbb{P}_{X,Z}\left(d(X,X^{\prime})\leq u\mid Z=Z^{\prime}\right),\,F_{Y^{\prime},Z^{\prime}}^{Y}(v)=\mathbb{P}_{Y,Z}\left(d(Y,Y^{\prime})\leq v\mid Z=Z^{\prime}\right).$
 We use $\Delta_{XY}(z)$ as measure of conditional profile association between $X$ and $Y$ given $Z=z$.
@@
 We provide a specific application example for the analysis of human longevity in Section 6.
-We focus here on the case where $Z$ is univariate and continuous.
-To avoid the curse of dimensionality under some structural assumptions, for the multivariate case one can apply single index Fréchet regression (Bhattacharjee and Müller, 2023), employing arguments that are similar to those in Zhou and Müller (2024a) for the case of multivariate predictors in conformal inference.
-For the univariate case we use nonparametric regression with argument $Z$ to implement conditional profiles, employing local linear estimators (Fan and Gijbels, 1992) for $F_{\omega_{1},z}^{X}(u)$, $F_{\omega_{2},z}^{Y}(v)$, and $F_{\omega_{1},\omega_{2},z}^{XY}(u,v)$.
+The definition (12) places no restrictions on the distribution of $Z$; hence, it remains valid whether $Z$ is continuous, discrete, or of mixed type.
+We focus on the case where $Z$ is univariate and continuous in the estimation procedure, as this setting is both commonly encountered in practice and theoretically nontrivial, particularly due to the challenges involved in establishing uniform convergence for local linear smoothing with object-valued data.
+When $Z$ is discrete (e.g., categorical), the estimation becomes significantly simpler: one can compute the association measure within each category of $Z$ and aggregate the results, avoiding the need for smoothing altogether.
+To avoid the curse of dimensionality under some structural assumptions, for the multivariate case one can apply single index Fréchet regression (Bhattacharjee and Müller, 2023), employing arguments that are similar to those in Zhou and Müller (2025) for the case of multivariate predictors in conformal inference.
+For the case of a univatiate continuous predictor $Z$ we use nonparametric regression with argument $Z$ to implement conditional profiles, employing local linear estimators (Fan and Gijbels, 1992) for $F_{\omega_{1},z}^{X}(u)$, $F_{\omega_{2},z}^{Y}(v)$, and $F_{\omega_{1},\omega_{2},z}^{XY}(u,v)$.
 This yields
@@
 $\hat{F}^{XY}_{\omega_{1},\omega_{2},z}(u,v)=$ $\mathrel{\mathop{\arg\min}\limits_{\beta_{0}\in\mathbb{R}}}\frac{1}{nh_{n}}\sum_{i=1}^{n}\left\{\mathds{1}_{\{d(X_{i},\omega_{1})\leq u\}}\mathds{1}_{\{d(Y_{i},\omega_{2})\leq v\}}-\beta_{0}-\beta_{1}(Z_{i}-z)\right\}^{2}$
-×$\quad\quad\quad\quad\times\mathrm{K}\left(\frac{Z_{i}-z}{h_{n}}\right),\hskip 369.88582pt$
-where $\mathrm{K}(\cdot)$ is a symmetric and Lipschitz continuous density kernel on $[-1,1]$ and $h_{n}$ is a sequence of bandwidths.
+$\quad\quad\quad\quad\times\mathrm{K}\left(\frac{Z_{i}-z}{h_{n}}\right),\hskip 369.88582pt$
+where $\mathrm{K}(\cdot)$ is a symmetric and Lipschitz continuous density kernel on [1,1][-1,1] and $h_{n}$ is a sequence of bandwidths.
 To study the asymptotic behavior of these conditional distance profile estimators, we introduce the sets of functions
-$\mathcal{F}_{1}$ =$=\,\{f(x)=\mathds{1}_{\{d(\omega_{1},x)\leq t\}}:\omega_{1}\in\mathcal{M}_{1}\,\,\,t\in{\mathbb{R}}^{+}\},$
-$\mathcal{F}_{2}$ =$=\,\{f(y)=\mathds{1}_{\{d(\omega_{2},y)\leq t\}}:\omega_{2}\in\mathcal{M}_{2},\,\,\,t\in{\mathbb{R}}^{+}\},$
+$\mathcal{F}_{1}$ $=\,\{f(x)=\mathds{1}_{\{d(\omega_{1},x)\leq t\}}:\omega_{1}\in\mathcal{M}_{1}\,\,\,t\in{\mathbb{R}}^{+}\},$
+$\mathcal{F}_{2}$ $=\,\{f(y)=\mathds{1}_{\{d(\omega_{2},y)\leq t\}}:\omega_{2}\in\mathcal{M}_{2},\,\,\,t\in{\mathbb{R}}^{+}\},$
 which are classes of indicator functions indexed by $\omega_{1}$, $\omega_{2}$ and $t$.
@@
 For every $\omega_{1}\in\mathcal{M}_{1}$, $\omega_{2}\in\mathcal{M}_{2}$ and $u,v\in{\mathbb{R}}^{+}$, $F^{XY}_{\omega_{1},\omega_{2},z}(u,v)$ is second order differentiable and has bounded second order derivatives with respect to $z$.
-Moreover, there exists a constant $L^{\prime}$ such that $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1}^{\prime},\omega_{2},z}(u,v)|\leq L^{\prime}d(\omega_{1},\omega_{1}^{\prime})$, $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1},\omega_{2}^{\prime},z}(u,v)|\leq L^{\prime}d(\omega_{2},\omega_{2}^{\prime})$ for all $z\in\mathcal{X},u,v\in{\mathbb{R}}^{+}$ and $\omega_{1},\omega_{1}^{\prime},\in\mathcal{M}_{1}$, $\omega_{2},\omega_{2}^{\prime},\in\mathcal{M}_{2}$.
+Moreover, there exists a constant $L^{\prime}$ such that $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1}^{\prime},\omega_{2},z}(u,v)|\leq L^{\prime}d_{1}(\omega_{1},\omega_{1}^{\prime})$, $|F^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1},\omega_{2}^{\prime},z}(u,v)|\leq L^{\prime}d_{2}(\omega_{2},\omega_{2}^{\prime})$ for all $z\in\mathcal{X},u,v\in{\mathbb{R}}^{+}$ and $\omega_{1},\omega_{1}^{\prime},\in\mathcal{M}_{1}$, $\omega_{2},\omega_{2}^{\prime},\in\mathcal{M}_{2}$.
 Proposition 2.
-Under assumptions 1 and 2, if $\max\{\mathcal{N}(\epsilon,\mathcal{F}_{1}),\,\mathcal{N}(\epsilon,\mathcal{F}_{2})\}\lesssim\epsilon^{-v}$ for some $v>0$ and $h_{n}\asymp(n/\log n)^{-1/5}$, then:
-[…]
-[…]
-[…]
-The first two statements of Proposition 4 follow directly from Theorem 3 in Zhou and Müller (2024a).
-Under the assumption $\max\{\mathcal{N}(\epsilon,\mathcal{F}_{1}),\,\mathcal{N}(\epsilon,\mathcal{F}_{2})\}\lesssim\epsilon^{-v}$, Lemma A.1 in Einmahl and Mason (2000) and Lemma S.4 in Zhou and Müller (2024b) imply that the function class $\mathcal{F}_{12}=\left\{f(x)=\mathds{1}_{\{d(\omega_{1},x)\leq t_{1}\}}\mathds{1}_{\{d(\omega_{2},x)\leq t_{2}\}}:\omega_{1}\in\mathcal{M}_{1},\,\omega_{2}\in\mathcal{M}_{1},\,t_{1},t_{2}\in\mathbb{R}\right\}$ has a polynomial covering number.
-The third statement is a consequence of Lemma 1 and Theorem 3 in Zhou and Müller (2024a).
+Under assumptions 1 and 4, if $\max\{\mathcal{N}(\epsilon,\mathcal{F}_{1}),\,\mathcal{N}(\epsilon,\mathcal{F}_{2})\}\lesssim\epsilon^{-v}$ for some $v>0$ and $h_{n}\asymp(n/\log n)^{-1/5}$,
+$\sup_{\omega_{1}\in\mathcal{M}_{1}}\sup_{z\in\mathcal{Z}}\sup_{u\geq 0}\left|\hat{F}^{X}_{\omega_{1},z}(u)-F^{X}_{\omega_{1},z}(u)\right|=O\left(\left(\frac{\log n}{n}\right)^{\frac{2}{5}}\right)\text{ a.s.},$
+$\sup_{\omega_{2}\in\mathcal{M}_{2}}\sup_{z\in\mathcal{Z}}\sup_{v\geq 0}\left|\hat{F}^{Y}_{\omega_{2},z}(v)-F^{Y}_{\omega_{2},z}(v)\right|=O\left(\left(\frac{\log n}{n}\right)^{\frac{2}{5}}\right)\text{ a.s.},$
+$\sup_{\omega_{1}\in\mathcal{M}_{1}}\sup_{\omega_{2}\in\mathcal{M}_{2}}\sup_{z\in\mathcal{Z}}\sup_{u,v\geq 0}\left|\hat{F}^{XY}_{\omega_{1},\omega_{2},z}(u,v)-F^{XY}_{\omega_{1},\omega_{2},z}(u,v)\right|=O\left(\left(\frac{\log n}{n}\right)^{\frac{2}{5}}\right)\text{ a.s.}.$
 An estimator for the conditional profile association ${\Delta}_{XY}(z)$ is then obtained through another nonparametric regression step,
@@
 Under the assumptions of Proposition 2,
-[…]
+$\sup_{z\in\mathcal{Z}}|\hat{\Delta}_{XY}(z)-{\Delta}_{XY}(z)|=O\left(h^{2}+\sqrt{\frac{\log n}{nh_{n}}}\right)\text{ a.s.}.$
 This result provides theoretical justification for the estimator (13) and supports its use as a measure for the conditional association of random effects.
@@
 $\mathcal{C}_{x,y,z}(u_{1},u_{2},v_{1},v_{2})\,=\,$ $C^{XY}_{x,y,z}(u_{1},u_{2},v_{1},v_{2})+F^{X}_{x,z}(u_{1})F^{X}_{x,z}(u_{2})C^{Y}_{y,z}(v_{1},v_{2})$ (15)
-$+F^{Y}_{y,z}(v_{1})F^{Y}_{y,z}(v_{2})C^{X}_{x,z}(u_{1},v_{2})+\mathcal{C}_{1}(u_{1},u_{2},v_{1},v_{2})+\mathcal{C}_{2}(u_{1},u_{2},v_{1},v_{2}),$
+$\hskip-85.35826pt+F^{Y}_{y,z}(v_{1})F^{Y}_{y,z}(v_{2})C^{X}_{x,z}(u_{1},u_{2})+\mathcal{C}_{1}(u_{1},u_{2},v_{1},v_{2})+\mathcal{C}_{2}(u_{1},u_{2},v_{1},v_{2}),$
 where
@@
 $p_{c}:=\frac{1+\sum_{\Pi_{j},\sigma_{jk}}\mathds{1}_{\{\lfloor n/2\rfloor\mathcal{T}_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XYZ}}(\Pi_{j},\sigma_{jk}))\geq n\mathcal{T}_{n}(\mathcal{D}_{\mathsf{XYZ}})\}}}{1+N},$
-where $N$ denotes the total number of permutations. The randomized test for testing (11) at significance level $\alpha$ is then defined as $\psi_{\alpha}^{c}:=\mathbb{P}(p_{c}\leq\alpha)$.
-Denoting by $\Gamma(z)$ the cumulative distribution function the limiting distribution in Theorem 6, its corresponding permutation-based estimator is
+where $N$ denotes the total number of permutations. The randomized test for testing (11) at significance level $\alpha$ is then defined as $\psi_{\alpha}^{c}:=\mathds{1}(p_{c}\leq\alpha)$.
+Denoting by $\Gamma(z)$ the cumulative distribution function of the limiting distribution in Theorem 6, its corresponding permutation-based estimator is
 $\hat{\Gamma}_{N}(z)=\frac{1}{N}\sum_{(\Pi_{j},\sigma_{jk})}\mathds{1}_{\left\{\frac{n}{2}\mathcal{T}_{\lfloor n/2\rfloor}(\mathcal{D}_{\mathsf{XYZ}}(\Pi_{j},\sigma_{jk}))\leq z\right\}},$
@@
 Corollary 1.
-Under the same assumptions as in Theorem 6, $|\hat{\Gamma}_{N}(z)-\Gamma(z)|=o_{P}(1)$ holds for every continuity point $z$ of $\Gamma(z)$, and $\mathbb{P}(\psi_{\alpha}^{c})\leq\alpha$.
+Under the same assumptions as in Theorem 6, $|\hat{\Gamma}_{N}(z)-\Gamma(z)|=o_{P}(1)$ holds for every continuity point $z$ of $\Gamma(z)$.
 5 Simulations
-We report here finite sample comparisons of the performance of the proposed profile approach with various alternative methods when $X$ and $Y$ are located in pertinent metric spaces.
-These include the Euclidean space $\mathbb{R}^{p}$, the sphere $\mathbb{S}^{2}$ and the 2-Wasserstein space of one-dimensional distributions $\mathbb{W}^{2}$ with the Wasserstein metric, where the latter is
-$d_{W}(x,y)^{2}=\int_{0}^{1}\{Q_{x}(u)-Q_{y}(u)\}^{2}\,{\rm d}u$ (16)
-and $Q_{x},\,Q_{y}$ are the quantile functions associated with the distributions $x$ and $y$.
-For each setting, both the proposed and comparison methods were implemented using the permutation test described in Section 3.2 and empirical power curves at significance level 0.05 are obtained from 500 Monte Carlo runs.
-We consider the following settings, where in all settings the dependence is quantified by the parameter $\rho$, where $\rho=0$ corresponds to independence and the level of dependence (association) increases with $\rho$.
-We obtained empirical power curves that illustrate how the power of the profile independence test increases as the argument $\rho$ increases, where we consider $0\leq\rho\leq 1$
+We present finite-sample comparisons of the proposed test against several alternative methods when $X$ and $Y$ are located in different metric spaces.
+These include the Euclidean space $\mathbb{R}^{p}$, the sphere $\mathbb{S}^{p}$, the space of $d$-dimensional symmetric positive definite matrices $\mathcal{S}_{p}^{++}$, and the 2-Wasserstein space of one-dimensional distributions.
+For the Euclidean space $\mathbb{R}^{p}$ and the sphere $\mathbb{S}^{p}$, we use the usual $\mathcal{L}^{2}$ distance and the geodesic metric, respectively.
+For the $\mathcal{S}_{p}^{++}$ space, we use the affine-invariant Riemannian distance $d(A,B)=\|\mathfrak{log}(A^{-1}B)\|_{F}^{2}$, where $\mathfrak{log}$ and $\mathfrak{exp}$ are the logarithm and exponential functions for matrices.
+Moreover, for $A,B\in\mathcal{S}_{p}^{++}$, the affine-invariant geodesic interpolation is defined by
+$A\#_{\rho}B=A^{1/2}(A^{-1/2}BA^{-1/2})^{\rho}A^{1/2}.$
+For the Wasserstein space $\mathbb{W}^{2}$, we use the 2-Wasserstein metric, which is defined as
+$d_{W}(x,y)^{2}=\int_{0}^{1}\{Q_{x}(u)-Q_{y}(u)\}^{2}\,{\rm d}u.$ (16)
+Here $Q_{x},\,Q_{y}$ are the quantile functions associated with the distributions $x$ and $y$.
+For each setting, both the proposed and comparison methods were implemented using the permutation-based test and empirical power curves at significance level 0.05 are obtained from 400 Monte Carlo runs.
+For the unconditional testing problem in Section 3, we consider the following settings, where in all settings the dependence is quantified by the parameter $\rho$, where $\rho=0$ corresponds to independence and the level of dependence (association) increases with $\rho$.
+We obtained empirical power curves that illustrate how the power of the profile independence test increases as the argument $\rho$ increases, where we consider $0\leq\rho\leq 1.$
+  • $\mathbb{R}^{p}$-Lin: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are related only through their first $J_{p}=\max\{1,\lfloor 0.1p\rfloor\}$ components, where each $Y_{j}$ for $j\leq J_{p}$ is generated as $Y_{j}=\rho X_{j}+\varepsilon_{j},$ with $\varepsilon_{j}\sim\mathcal{N}(0,1)$ independent of all other variables.
+The remaining components $Y_{J_{p}+1},\ldots,Y_{p}$, and $X_{1},\ldots,X_{p}$, are independently distributed as $\mathcal{N}(0,1)$.
+  • $\mathbb{R}^{p}$-Log: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are generated as in the $\mathbb{R}^{p}$-Lin setting, except that $Y_{j}=\rho\log(4X_{j}^{2})+\varepsilon_{j}$, where $\varepsilon_{j}\sim\mathcal{N}(0,1)$ is independent of all other variables.
   • $\mathbb{R}^{1}$-Cir: Denoting by $\|v\|_{\zeta}$ the $\zeta$-norm for any Euclidean vector $v$.
 For any $1\leq\zeta\leq\infty$ we consider pairs $(X,Y)$ of one-dimensional random variables $X$ and $Y$ that satisfy the condition $\|(X,Y)\|_{2/\rho}\leq 1$, where $X$ and $Y$ are independently drawn from $\text{Unif}(-1,1)$.
-When $\rho=0$, $(X,Y)$ is uniformly distributed over the square $[-1,1]\times[-1,1]$.
-When $\rho=1$, $(X,Y)$ follows a uniform distribution within the unit circle $\{(x,y)\mid x^{2}+y^{2}\leq 1\}$, with domains for the uniform distribution located in between the square and the circle for $10<\rho<$ 1 .
-  • $\mathbb{R}^{1}$-Bimodal: The pair $(X,Y)$ of one-dimensional random variables satisfies $Y=\epsilon|X|^{\rho}$, where $X$ and $\epsilon$ are independent and follow mixture distributions $0.2\mathcal{N}(-3,1)+0.8\mathcal{N}(3,1)$.
-  • $\mathbb{R}^{p}$-Log: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are related only through their first elements, specifically $Y_{1}=\rho\log(4X_{1}^{2})+\epsilon$.
-The variables $Y_{2},\ldots,Y_{p}$, $X_{1},X_{2},\ldots,X_{p}$, and $\epsilon$ are all distributed as $\mathcal{N}(0,1)$ and are independent.
-  • $\mathbb{R}^{p}$-Pow: $\bm{X}=(X_{1},\ldots,X_{p})^{\top}$ and $\bm{Y}=(Y_{1},\ldots,Y_{p})^{\top}$ are related only through the first element, where $Y_{1}=\epsilon\times(|X_{1}|^{\rho})$.
-The variables $Y_{2},\ldots,Y_{p}$ and $\epsilon$ are distributed as $\mathcal{N}(0,1)$, while $X_{1},X_{2},\ldots,X_{p}$ are distributed as $\mathrm{Unif}(-1,1)$.
-Additionally, $Y_{2},\ldots,Y_{p}$, $X_{1},X_{2},\ldots,X_{p}$, and $\epsilon$ are independent.
-  • $\mathbb{S}^{2}$-Lin: $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by independent standard normal random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$:
-    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
-    $Y_{1}=\frac{\rho\epsilon_{1}+(1-\rho)\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|},$
-    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(\rho\epsilon_{1}+(1-\rho)\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
-  • $\mathbb{S}^{2}$-Log: $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by independent standard normal random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$:
-    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
-    $Y_{1}=\frac{\rho\log(4\epsilon_{1}^{2})+\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|},$
-    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(\rho\log(4\epsilon_{1}^{2})+\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
-  • $\mathbb{S}^{2}$-Pow: $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by independent standard normal random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$:
-    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
-    $Y_{1}=\frac{|\epsilon_{1}|^{\rho}+\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|},$
-    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(|\epsilon_{1}|^{\rho}+\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
-  • $\mathbb{S}^{2}$-Pow-Bimodal: Here $\bm{X}=(X_{1},X_{2},X_{3})$ and $\bm{Y}=(Y_{1},Y_{2},Y_{3})$ are random elements on the unit sphere $\mathbb{S}^{2}$, generated by the random variables $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{6}$, where $\epsilon_{i}$ are random samples from a mixture of Gaussians $0.2\mathcal{N}(-1,1)+0.8\mathcal{N}(1,1)$:
-    $X_{1}=\frac{\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\,X_{3}=\frac{\epsilon_{3}}{\|\bm{\epsilon}_{X}\|};$
-    $Y_{1}=\frac{|\epsilon_{1}|^{\rho}+\epsilon_{4}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{5}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{3}=\frac{\epsilon_{6}}{\|\bm{\epsilon}_{Y}\|}$
-    where $\bm{\epsilon}_{X}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})^{\top}$ and $\bm{\epsilon}_{Y}=(|\epsilon_{1}|^{\rho}+\epsilon_{4},\epsilon_{5},\epsilon_{6})^{\top}$.
-  • $\mathbb{W}^{2}$-Mean: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(\mu_{X},1)$ and $Y=\mathcal{N}(\mu_{Y},1)$, where $\mu_{X}\sim\mathrm{Unif}(-1,1)$, $\mu_{Y}=\epsilon\times|\mu_{X}|^{\rho}$, and $\epsilon\sim\mathcal{N}(0,1)$ and is independent of $\mu_{X}$.
-  • $\mathbb{W}^{2}$-Var: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(0,\sigma_{X}^{2})$ and $Y=\mathcal{N}(0,\sigma_{Y}^{2})$, where $\sigma_{X}^{2}\sim\mathrm{Unif}(0,1)$ and $\sigma_{Y}=\rho\sin(\sigma_{X}\pi)+\epsilon$, and where $\epsilon\sim\mathrm{Unif}(0,2)$ and is independent of $\sigma_{X}$.
-For cases where $X$ and $Y$ are in $\mathbb{R}^{p}$ for $p\geq 1$, we compare the proposed profile independence test with the energy test (Székely et al., 2007), the ball divergence test (Pan et al., 2020) and Chatterjee’s rank test (Chatterjee, 2021).
-All methods are implemented using R packages: the energy test (Energy) via dcov.test() from the energy package, the ball divergence test (Ball) via bcov.test() through the Ball package (Zhu et al., 2021) and Chatterjee’s rank test (Chatterjee) via xicor() through the XICOR package.
-To compare different approaches, we focus on the power curves as function of $\rho$ for the simulation settings described above.
-Figure 1 indicates that the proposed method outperforms all comparable methods, particularly in the $\mathbb{R}^{1}$-Bimodal setting, where the energy test and Chatterjee’s rank test fail.
-Additionally, in this more complex setting, the proposed test exhibits a faster increase in power as the sample size increases.
-Figure 2 shows that the proposed profile independence test (blue) continues to consistently outperform the energy test (black) and the ball divergence test (red) across all simulation settings where $X$ and $Y$ reside in multivariate Euclidean spaces or metric spaces.
-The simulations thus provide evidence that the profile independence test effectively adapts to complex dependency structures where other methods fail.
-For example, when the relationship between $X$ and $Y$ is linear, a scenario where also classical methods perform well, the energy test performs comparably to the profile independence test when the relationship between $X$ and $Y$ is linear.
-However, in more complex cases, such as log- or power-based dependencies, and especially in settings $\mathbb{S}^{2}$-Log and $\mathbb{S}^{2}$-Pow, the energy test performs poorly, whereas the profile independence test maintains strong detection power.
-These results provide evidence for the robustness and efficiency of the proposed test in detecting dependencies across diverse scenarios.
-Figure 1: Power curves comparing the proposed profile independence test (blue), the ball statistic test (red), the Chatterjee est (yellow) and the energy test (black), where power is a function of the increasing dependency parameter $\rho$, for two simulation settings: $\mathbb{R}^{1}$-Cir (first row) and $\mathbb{R}^{1}$-Bimodal (second row) The dashed grey line indicates the target significance level of $0.05$.
-Three different sample sizes are considered: $n=100$ (left column), $n=200$ (middle column), and $n=400$ (right column).
-Figure 2: Power curves comparing the proposed profile independence test (blue), the ball statistic test (red), the Chatterjee test (yellow) and the energy test (black), where power is a function of the increasing dependency parameter $\rho$, for nine simulation settings: $\mathbb{R}^{5}$-Log, $\mathbb{R}^{10}$-Log, and $\mathbb{R}^{10}$-Pow (left, middle, right panel in first row); $\mathbb{S}^{2}$-Lin, $\mathbb{S}^{2}$-Log, and $\mathbb{S}^{2}$-Pow (left, middle, right panel in second row); $\mathbb{S}^{2}$-Pow-Bimodal, $\mathbb{W}^{2}$-Mean, and $\mathbb{W}^{2}$-Var (left, middle, right panel in third row).
-The dashed grey line shows the target significance level of $0.05$, and the sample size is set to $n=400$.
+When $\rho=0$, $(X,Y)$ is uniformly distributed over the square $1,1]\times[-1,1]$.
+When $\rho=1$, $(X,Y)$ follows a uniform distribution within the unit circle $\{(x,y)\mid x^{2}+y^{2}\leq 1\}$, with domains for the uniform distribution located in between the square and the circle for $0<\rho<1.$
+  • $\mathcal{S}_{2}^{++}$-Interpolation: $X$ and $Y$ are $2\times 2$ symmetric positive definite matrices.
+$X$ is generated as $X=\mathfrak{exp}((S+S^{\top})/2)$, where $S$ is a $2\times 2$ matrix with independent entries drawn from $\mathcal{N}(0,0.6^{2})$.
+Let $Y_{(0)}$ denote the independent baseline, which is generated independently but follows the same distribution as $X$.
+The dependent target $Y_{(1)}$ incorporates geodesic noise with $P=CXC^{\top}$, and is constructed as
+    $Y_{(1)}=P^{1/2}\mathfrak{exp}(0.1(E+E^{\top})/2)P^{1/2},$
+    where $C={\rm diag}(2,0.5)$ and $E$ is a $2\times 2$ matrix with independent standard normal entries.
+The final matrix $Y$ is obtained via geodesic interpolation between the independent baseline $Y_{(0)}$ and the dependent target $Y_{(1)}$, defined as
+    $Y=Y_{(0)}\#_{\rho/5}Y_{(1)}=Y_{(0)}^{1/2}(Y_{(0)}^{-1/2}Y_{(1)}Y_{(0)}^{-1/2})^{\rho/5}Y_{(0)}^{1/2}.$
+  • $\mathbb{S}^{2}$$\mathcal{S}_{2}^{++}$-Hybrid: Here we consider $X\in\mathbb{S}^{2}$ as on the unit sphere, while $Y\in\mathcal{S}_{2}^{++}$ is a symmetric positive definite matrix.
+First, generate $X_{(0)}$ and $Y$ as in the setting $\mathcal{S}_{2}^{++}$-Interpolation.
+Then, $X$ is the vectorized form of the upper triangular part of $X_{(0)}$, normalized so that $\|X\|_{2}=1$.
+  • $\mathbb{W}^{2}$-Mean: $X$ and $Y$ are random distributions in $\mathbb{W}^{2}$, specifically $X=\mathcal{N}(\mu_{X},\sigma_{X})$ and $Y=\mathcal{N}(\mu_{Y},\sigma_{Y})$, where $\mu_{X}\sim\mathcal{N}(0,1)$, $\mu_{Y}=0.2\epsilon\times|\mu_{X}|^{\rho}$, $\sigma_{X}\sim\mathrm{Unif}(0,1)$, $\sigma_{Y}\sim\mathrm{Unif}(1,2)$ and $\epsilon\sim\mathcal{N}(0,1)$.
+We compare the proposed profile-based independence test with the energy test (Székely et al., 2007) and the ball divergence test (Pan et al., 2020).
+All methods are implemented using available R packages: the energy test via dcov.test() from the energy package, the ball divergence test via bcov.test() from the Ball package (Zhu et al., 2021), and Chatterjee’s rank test via xicor() through the XICOR package.
+Since Chatterjee’s rank test only supports scalar independence testing, we include it only for the $\mathbb{R}^{1}$ setting.
+Figure 1 presents power curves as functions of the dependency parameter $\rho$ for the simulation settings described above.
+The simulation results show that the proposed profile-based test effectively adapts to complex dependency structures where traditional methods tend to fail.
+For instance, in the Euclidean space $\mathbb{R}^{2}$, when $Y$ and $X$ exhibit linear dependence, the energy test slightly outperforms the profile-based test.
+However, the proposed method performs better when the dependency becomes logarithmic.
+In the more complex setting $\mathbb{R}^{1}$-Cir, both the energy test and Chatterjee test fail, while the proposed test achieves good power and outperforms the ball divergence test.
+Moreover, in more challenging scenarios involving non-Euclidean spaces $\mathcal{S}_{2}^{++}$, $\mathbb{S}^{2}$, and $\mathbb{W}^{2}$, the proposed test performs significantly better than all other comparable methods.
+These results are consistent with our theoretical findings and highlight the robustness and efficiency of the proposed test across a wide range of dependency structures.
+In summary, the proposed method performs comparably to existing approaches under linear dependence and outperforms them when the dependency structure is more intricate.
+Figure 1: Power curves comparing the proposed profile-based independence test (blue solid lines), the ball divergence test (black dashed lines), the energy test (red dotted lines) and the Chatterjee test (yellow dash-dotted lines), where power is plotted as a function of the dependency parameter $\rho$, defined in the second paragraph of section 5, across six simulation settings: $\mathbb{R}^{1}$-Cir, $\mathbb{R}^{2}$-Log, and $\mathbb{R}^{10}$-Log (left, middle, and right panels in the first row); $\mathcal{S}_{2}^{++}$-Interpolation, $\mathbb{S}^{2}$$\mathcal{S}_{2}^{++}$-Hybrid, $\mathbb{W}^{2}$-Mean (left, middle, and right panels in the second row).
+The dashed grey line indicates the target significance level of 0.05, and the sample size is set to $n=200$.
+We next study the performance of the proposed method in several high-dimensional settings.
+Figure 2 presents the rejection rates as a function of sample size for the settings $\mathbb{R}^{100}$-Log and $\mathbb{W}^{2}$-Mean under both $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$.
+The results demonstrate that the proposed method consistently controls the type I error across all sample sizes.
+In the Euclidean $\mathbb{R}^{100}$-Log setting, the proposed method outperforms the energy test and performs comparably to the ball divergence test, with slightly better power for large sample sizes.
+This suggests that, although the proposed method is not specifically tailored for high-dimensional data, it still achieves competitive or superior finite-sample performance.
+In the infinite-dimensional $\mathbb{W}^{2}$-Mean setting, the proposed method exhibits a near-perfect power curve for relatively large sample sizes, while all other methods fail and show limited rejection rates.
+We include results for additional settings in the Supplement, illustrating how the power curves evolve with increasing sample size and approach perfect power for large samples.
+Additionally, we assess the performance of the proposed test across a range of significance levels in the Supplement.
+Figure 2: Rejection rate of the proposed independence test for two high-dimensional settings (left two panels: $\mathbb{R}^{100}$-Log, right two panels: $\mathbb{W}^{2}$-Mean).
+Different colors and linetypes represent different methods: blue solid for the proposed method, black dashed for the ball divergence test, and red dotted for the energy test.
+Within each setting, two separate plots show the results under $\mathcal{H}_{0}$ (left) and $\mathcal{H}_{1}$ (right).
+The dashed grey line indicates the target significance level of 0.05
+Figure 3: Power curves of the proposed conditional independence test across four simulation settings with different metric spaces: Conditional-$\mathbb{S}^{2}$-Log, Conditional-$\mathbb{S}^{4}$-Log, Conditional-$\mathbb{W}^{2}$-Log, and Conditional-$\mathbb{W}^{2}$-Sin.
+Each plot shows the rejection rate as a function of the dependency parameter $\rho$, defined in the second paragraph of section 5, with sample size fixed at $n=200$.
+The covariate $Z$ is drawn from $\mathrm{Unif}[0,1]$, and each point is based on 500 Monte Carlo replications.
+The dashed grey line indicates the target significance level of 0.05.
+For the conditional independence testing, since there are no existing methods designed specifically for random objects, we focus on evaluating the performance of the proposed method in several metric space settings.
+In all simulation settings below, the covariate $Z$ is generated from a $\mathrm{Unif}[0,1]$ distribution, and $\rho$ denotes the strength of dependence as defined in the second paragraph of section 5,
+  • Conditional-$\mathbb{S}^{p}$-Log: $\bm{X}=(X_{1},X_{2},\ldots,X_{p+1})$ and $\bm{Y}=(Y_{1},Y_{2},\ldots,Y_{p+1})$ are random elements on the unit sphere $\mathbb{S}^{p}$, generated with $Z$ and independent standard normal random variable $\epsilon_{1},\,\epsilon_{2},\ldots,\epsilon_{2p+1}$:
+    $X_{1}=\frac{Z+\epsilon_{1}}{\|\bm{\epsilon}_{X}\|},\,\,X_{2}=\frac{\epsilon_{2}}{\|\bm{\epsilon}_{X}\|},\,\ldots,\,X_{p+1}=\frac{\epsilon_{p+1}}{\|\bm{\epsilon}_{X}\|};$
+    $Y_{1}=\frac{Z+\rho\log(4\epsilon_{1}^{2})+(1-\rho)\epsilon_{p+2}}{\|\bm{\epsilon}_{Y}\|},\,\,Y_{2}=\frac{\epsilon_{p+3}}{\|\bm{\epsilon}_{Y}\|},\,\ldots,\,\,Y_{3}=\frac{\epsilon_{2p+2}}{\|\bm{\epsilon}_{Y}\|},$
+    where $\bm{\epsilon}_{X}=(Z+\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{p+1})^{\top}$, $\bm{\epsilon}_{Y}=(Z+\rho\log(4\epsilon_{1}^{2})+(1-\rho)\epsilon_{p+2},\epsilon_{p+3},\ldots,\epsilon_{2p+2})^{\top}$.
+  • Conditional-$\mathbb{W}^{2}$-Log: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(\mu_{X},1)$ and $Y=\mathcal{N}(\mu_{Y},1)$, where $\mu_{X}\sim Z+\mathrm{Unif}(-1,1)$, $\mu_{Y}=Z+\rho\log(4\mu_{X}^{2})+\epsilon$, and $\epsilon\sim\mathrm{Unif}(-1,1)$ and is independent of $\mu_{X}$ and $Z$.
+  • Conditional-$\mathbb{W}^{2}$-Sin: $X$ and $Y$ are random elements in the space of distributions $\mathbb{W}^{2}$ with the Wasserstein metric, specifically $X=\mathcal{N}(\mu_{X},1)$ and $Y=\mathcal{N}(\mu_{Y},1)$, where $\mu_{X}\sim Z+\mathrm{Unif}(-1,1)$, $\mu_{Y}=Z+\rho\sin(\mu_{X}\pi)+\epsilon$, and $\epsilon\sim\mathrm{Unif}(-1,1)$ and is independent of $\mu_{X}$ and $Z$.
+The results are shown in Figure 3.
+The proposed conditional test shows increasing power as the dependency parameter $\rho$ grows, achieving near-perfect rejection rates for moderate to strong dependence.
+These results highlight the adaptability and robustness of the proposed method across different metric spaces and across complex nonlinear scenarios such as the log and sinusoidal dependence structures.
 6 Data illustrations
@@
 The data are publicly accessible at https://www.humanconnectome.org.
-For our analysis, we focus on a subset comprising $n=209$ young, healthy adults who completed both resting-state and motor task fMRI sessions.
-The resting-state fMRI data were collected while participants were instructed to keep their eyes open and fixate on a projected bright cross-hair on a dark background in a darkened room.
-For the motor task fMRI recordings, participants were presented with visual cues prompting them to perform specific movements, such as squeezing their left or right toes or moving their tongue, to map specific motor areas.
+For our analysis, we focus on a subset comprising $n=209$ young healthy adults who completed both resting-state and motor task fMRI sessions.
+The resting-state fMRI data were collected while participants were instructed to keep their eyes open and fixate on a projected bright cross-hair on a dark background.
+For the motor task fMRI recordings, participants were presented with visual cues prompting them to perform specific movements, such as squeezing their left or right toes or moving their tongue.
 Each time block for the recording of a specific movement type lasted 12 seconds, included 10 movements and was preceded by a 3-second cue.
@@
 For instance, the correlation matrix for “squeezing left toe” was derived from the time periods during which participants were actively performing the “squeezing left toe” task.
-For each subject, this resulted in a total of eight $9\times 9$ correlation matrices, corresponding to one rest phase and three task phases (resting state, squeezing left toe, squeezing right toe, and moving tongue) for each of the two brain hemispheres (left and right).
-Figure 3 shows the estimated Fréchet mean of these eight covariance matrices when employing the Frobenius norm.
-This indicates that the task-evoked correlation matrices display higher correlations between different brain regions compared with the resting-state correlation matrices.
-Figure 3: Fréchet mean correlation matrices for resting-state fMRI (left-most column) and task-evoked fMRI (next three columns: squeezing left toe, squeezing right toe, and moving tongue).
-The top row represents the left hemisphere and the bottom row the right hemisphere.
-We then proceeded to apply the proposed profile independence test, along with the energy test (Székely et al., 2007), aiming to test the independence of random object pairs $\{(X_{i},Y_{i})\}_{i=1}^{209}$.
-Here $X_{i}$ represents the correlation matrix corresponding to resting-state fMRI, and $Y_{i}$ the correlation matrix corresponding to task-evoked fMRI, for the various motor tasks for the $i$th subject.
+For each subject, this resulted in a total of four $18\times 18$ correlation matrices for the whole brain, corresponding to one rest phase and three task phases (resting state, squeezing left toe, squeezing right toe, and moving tongue).
+Figure 4 shows the estimated Fréchet mean of the four covariance matrices under the Frobenius norm.
+The results indicate that task-evoked correlation matrices exhibit stronger correlations across different brain regions compared to resting-state correlation matrices.
+Furthermore, in both resting-state and task-evoked conditions, the connectivity matrices between the left and right hemispheres are highly correlated.
+Applying both the proposed independence test and the energy test to these hemispheric connectivity matrices yields $p$-values below 0.01 in all cases, confirming that the left and right brain connectivities are not independent, which is consistent with expectations.
+Figure 4: Fréchet mean correlation matrices for resting-state fMRI (leftmost column) and task-evoked fMRI (next three columns: squeezing the left toe, squeezing the right toe, and moving the tongue).
+In each matrix, the top-left $9\times 9$ block represents the connectivity within the left hemisphere, and the bottom-right $9\times 9$ block represents the connectivity within the right hemisphere.
+Given the strong dependence between left and right hemisphere connectivity matrices observed across both resting-state and task-evoked conditions, we focus our analysis on separate $9\times 9$ intra-hemispheric correlation matrices.
+This is motivated by the fact that when left and right hemispheres are highly correlated, performing independence tests on the full $18\times 18$ whole-brain matrices can make it easier to reject the null hypothesis of independence due to inflated signal from inter-hemispheric correlations.
+In contrast, tests based on the smaller $9\times 9$ matrices are more stringent and informative for evaluating within-hemisphere dynamics.
+To this end, we applied the proposed test and the energy test (Székely et al., 2007) to assess the independence of random object pairs $\{(X_{i},Y_{i})\}_{i=1}^{209}$.
+Here, $X_{i}$ denotes the correlation matrix from the resting-state fMRI, and $Y_{i}$ denotes the correlation matrix from the task-evoked fMRI corresponding to the various motor tasks, for the left and right hemispheres of the $i$th subject.
 We consider three metrics for the space of $9\times 9$ positive definite matrices $A$ and $B$.
-Specifically, we compare the results for the following three metrics that were applied previously in the literature for various data analysis tasks:
+Specifically, we compare the results for the following metrics that were applied previously in the literature for various data analysis tasks:
   • a)
@@
 Here $\lfloor L_{A}\rfloor$ denotes the off-diagonal part of $L_{A}$ and $L_{A,jj}$ the $j$th diagonal element of $L_{A}$.
+  • c)
+    $\alpha$-Power metric (Dryden et al., 2009; Zhou and Mülller, 2022)
+    $d_{F,\alpha}=\|A^{\alpha}-B^{\alpha}\|_{F},$
+    where $A^{\alpha}=U\Lambda^{\alpha}U^{\top}$ and $A=U\Lambda U^{\top}$ is the usual spectraL decomposition of a positive definite matrix AA.
+  • d)
+    Bures-Wasserstein metric
+    $d_{{BW}}(A,B)^{2}=\operatorname{Tr}(A+B-2(A^{1/2}BA^{1/2})^{1/2}).$
 For each metric and method, we conducted six tests in total: three for each motor task (“squeezing the left toe”, “squeezing the right toe”, and “moving the tongue”) separately for the left and right hemispheres.
 The $p$-values resulting from the permutation test implementation of the proposed profile independence test are shown in Table 1 and are also compared with the results from the energy test for independence.
-For the tongue movement task (columns 3, 6, and 9), both the proposed profile independence test and the energy test yield the same result across all three metrics: there is no evidence to reject the null hypothesis that the correlation matrices between the resting state and the tongue movement task are independent.
-For the toe-squeezing task, under the Frobenius metric (columns 1 and 2), both the proposed profile independence test and the energy test indicate dependence at the 0.05 significance level.
-However, the energy test fails to detect dependence under the AIRM and Log-Cholesky metrics (row 2, columns 5 and 8), whereas the proposed profile independence test consistently rejects the null hypothesis in all these cases.
+For the tongue movement task, both the proposed profile independence test and the energy test yield the same result across all three metrics: there is no evidence to reject the null hypothesis that the correlation matrices between the resting state and the tongue movement task are independent.
+For the toe-squeezing task, under the Frobenius metric, both the proposed profile independence test and the energy test indicate dependence at the 0.05 significance level.
+However, the energy test fails to detect dependence under the AIRM, Log-Cholesky, $1/2$-Power, $1/3$-Power and Bures–Wasserstein metrics for the right toe-squeezing task, whereas the proposed profile independence test consistently rejects the null hypothesis in all these cases.
 These results demonstrate the consistency of the proposed test across different metric choices.
+For completeness, we also include test results based on the full $18\times 18$ correlation matrices in the Supplement, with the above caveat.
+Both the proposed method and the energy test yield consistently low $p$-values for the left and right toe-squeezing tasks, whereas the $p$-values for the tongue movement task are comparatively higher.
+These results are consistent across all metrics.
 Table 1: Comparison of $p$-values obtained when applying the proposed profile independence test and the energy test for independence for the null hypothesis of independence between resting-state fMRI and task-evoked fMRI connectivity matrices for three motor tasks: $\mathrm{Ltoe}$ (squeezing the left toe), $\mathrm{Rtoe}$ (squeezing the right toe), and $\mathrm{tongue}$ (moving the tongue).
-The results are presented for left and right hemispheres and for three different metrics: Frobenius metric (left three columns), Affine-invariant Riemannian metric (middle three columns) and Log-Cholesky metric (right three columns).
-The first two rows correspond to correlation matrices show the results for the left hemisphere and the last two rows for the right hemisphere.
+The results are presented for left and right hemispheres and for five different metrics: Frobenius metric (left three columns, upper table), Affine-invariant Riemannian metric (middle three columns, upper table), Log-Cholesky metric (right three columns, upper table), $1/2$-Power metric (left three columns, lower table), $1/3$-Power metric (middle three columns, lower table) and Bures–Wasserstein metric (right three columns, lower table).
+The first two rows in upper/lower table show the results for the correlation matrices of the left hemisphere and the last two sub-rows for the right hemisphere.
 The $p$-values for the proposed profile independence test and the energy test were obtained through permutation tests with a total number of permutations $N=500$ and are bolded if $p>0.1$.
@@
 energy 0.00 0.00 0.13 0.00 0.00 0.20 0.02 0.02 0.22
+1/2-Power metric 1/3-Power metric Bures-Wasserstein
+Ltoe Rtoe tongue Ltoe Rtoe tongue Ltoe Rtoe tongue
+Left proposed 0.00 0.03 0.55 0.00 0.02 0.68 0.00 0.03 0.54
+energy 0.01 0.25 0.55 0.01 0.37 0.57 0.01 0.26 0.55
+Right proposed 0.02 0.03 0.42 0.01 0.02 0.49 0.02 0.03 0.40
+energy 0.00 0.00 0.22 0.01 0.00 0.25 0.01 0.00 0.22
 6.2 Age-at-death distributions
 The study of human mortality as quantified by age-at-death distributions has attracted much interest over the past decades.
-The analysis of age-at-death distributions has provided insights into mortality patterns, lifespan variability and the factors influencing survival across different populations and time periods.
+The analysis of age-at-death distributions has led to deeper insights into mortality patterns, lifespan variability and the factors influencing survival across different populations and time periods.
 The Human Mortality Database (www.mortality.org) provides annual age-at-death tables for 38 countries, with death rates recorded for each country across ages ranging from 0 to 110.
@@
 Here $X$ stands for the male age-at-death distribution for a given country at calendar year $Z$, and $Y$ the corresponding age-at-death distribution for females for the same country and year.
-We employ the Wasserstein-2 metric (16) in the space of one-dimensional distributions considered here, as it is particularly sensitive to horizontal shifts in distributions that are of special interest.
-We apply the proposed conditional profile association $\Delta_{XY}(z)$ in (12) for this analysis, aiming to shed light on the changing nature of the relationship between male and female age-at-death distributions across different calendar years.
-The conditional profile association serves to quantify the degree of dependence between the two distributions as a function of calendar time $z$, where a smaller value of $\Delta_{XY}(z)$ indicate that male and female death rate distributions are relatively less dependent.
-dependent, while larger values indicate stronger association and a higher degree of dependence.
-We divide the 33 countries into two groups: Eastern European countries, which include Bulgaria, Belarus, Czechia, Estonia, Hungary, Lithuania, Latvia, Poland, Slovakia, and Slovenia; and other countries that can be broadly characterized as Western/Asian and include Australia, Austria, Belgium, Canada, Switzerland, Denmark, Spain, Finland, France, England and Wales, Greece, Ireland, Iceland, Italy, Japan, Luxembourg, the Netherlands, Norway, New Zealand, Portugal, Sweden, Taiwan and the United States.
-The results are in Figure 4, which illustrates the estimated conditional profile association as a function of calendar year for all countries combined, as well as separately for Eastern European and Western/Asian countries.
+We use calendar time as covariate $Z$, an inherently continuous variable, although it is initially observed only at discrete time points, taken to be the midpoints of calendar years.
+With 35 consecutive years in the dataset and mortality rates exhibiting smooth temporal variation, this observation scheme allows for the effective application of local smoothing methods in the pre-processing.
+We employ the Wasserstein-2 metric (16) in the space of one-dimensional age-at-death distributions, as it is particularly sensitive to horizontal shifts in distributions that are of special interest.
+We applied the proposed conditional profile association $\Delta_{XY}(z)$ in (12) for this analysis, aiming to shed light on the changing nature of the relationship between male and female age-at-death distributions across different calendar years.
+The conditional profile association serves to quantify the degree of dependence between the two distributions as a function of calendar time $z$, where a smaller value of $\Delta_{XY}(z)$ indicates that male and female death rate distributions are relatively less dependent, and vice versa.
+We divided the 33 countries into two groups: Eastern European countries (Bulgaria, Belarus, Czechia, Estonia, Hungary, Lithuania, Latvia, Poland, Slovakia, and Slovenia) and Western/Asian countries (Australia, Austria, Belgium, Canada, Switzerland, Denmark, Spain, Finland, France, England and Wales, Greece, Ireland, Iceland, Italy, Japan, Luxembourg, the Netherlands, Norway, New Zealand, Portugal, Sweden, Taiwan and the United States).
+The results are in Figure 5, which illustrates the estimated conditional profile association as a function of calendar year for all countries combined and separately for Eastern European and Western/Asian countries.
 The left panel, representing all countries, exhibits a generally decreasing trend over time, indicating a gradual reduction in the association between male and female age-at-death distributions over the 25 years covered in this analysis.
@@
 The association in Western countries also follows a steady downward trend, suggesting an increasing independence between male and female mortality distributions over time.
-This lower and declining association seems to indicate an increasing decoupling between male and female mortality over time that is much more expressed for Western/Asian countries but is much less evident for Eastern European countries.
+This lower and declining association seems to indicate an increasing decoupling between male and female mortality over time that is more expressed for Western/Asian countries.
 The clear disparity between Western and Eastern European countries highlights the varying demographic, political and health-related forces that influence the relationship between male and female mortality.
+To complement the observed decreasing trend in conditional profile associations between male and female mortality, we conducted the proposed conditional independence test described in Section 4.
+The resulting $p$-values are consistently below 0.01 for both Eastern European and Western/Asian countries, providing strong statistical evidence against conditional independence over the time period considered.
+These findings are expected and confirm that mortality rates for males and females are overall dependent, as expected.
 Figure 5: Estimated conditional profile associations as a function of calendar year for distributional data that correspond to male and female age-at-death distributions across 33 countries.
@@
 This extension provides a useful tool for exploring relationships between complex object responses and Euclidean covariates, with immediate practical relevance.
-Beyond independence testing, the profile association framework has potential for broader statistical applications, including causal inference, change-point detection, and feature selection, thus opening promising avenues for future research.
+Beyond independence testing, the profile association framework has potential for broader statistical applications, including causal inference.
 Causal inference has emerged as an important field in statistics, with significant applications across various scientific disciplines.
-One key challenge in causal inference is detecting confounding contributions, which represent spurious associations arising from the influence of a confounder on two outcomes of interest.
+One key challenge in causal inference is detecting confounding contributions that can lead to spurious associations arising from the influence of a confounder on two outcomes of interest.
 A simple conceptual tool for understanding confounding for Euclidean data is the classical law of total covariance,
@@
 which decomposes the conditional covariance between random variables $X$ and $Y$ in the presence of a conditioning variable $Z$.
-Here $\text{Cov}(\mathbb{E}[X\mid Z],\mathbb{E}[Y\mid Z])$ is the contribution to the covariance explained by $Z$, and thus possibly due to a confounding effectof $Z$, whereas $\text{Cov}(\mathbb{E}[X\mid Z],\mathbb{E}[Y\mid Z])$ is the part that is not explained by $Z$ in the context of a linear framework.
-An interesting question for future research is whether by utilizing the proposed conditional profile association one can extend such a separation of the effect that can be attributed to $Z$ and the remaining association that cannot be explained by $Z$ to scenarios where $X$ and $Y$ are random objects and a linear framework does not apply, due to the lack of vector operations in non-Euclidean spaces.
-Considering the mortality example discussed in Section 6.2, where $(X_{i}$ and $Y_{i}$ are female and male age-at-death distributions, while $Z_{i}$ is a Euclidean covariate, a preliminary attempt to shed some light on this is the following three-step procedure: first compute the unconditional profile association $\Delta(X,Y)$, using all available data pairs $\{(X_{i},Y_{i})\}_{i}$; second, replace the conditional expectations $\mathbb{E}[X\mid Z]$, $\mathbb{E}[Y\mid Z])$ in the law of total covariance by Fréchet regressions (Petersen and Müller, 2019) to obtain conditional Fréchet means $\mathbb{E}_{\oplus}[X\mid Z]$ and $\mathbb{E}_{\oplus}[Y\mid Z]$ using data $\{(X_{i},Z_{i})\}_{i}$ and $\{(Y_{i},Z_{i})\}_{i}$ respectively; third, compute $\Delta(\mathbb{E}_{\oplus}[X\mid Z],\mathbb{E}_{\oplus}[Y\mid Z])$ by plugging in the results from the Fréchet regression step.
-Applying these considerations for the human mortality data for all countries combined yields $0.3630\Delta(X,Y)=$ 0.36 for the estimate of the overall profile association that ignores calendar year, but $30\Delta(\mathbb{E}[_{\oplus}X\mid Z],\mathbb{E}[_{\oplus}Y\mid Z])=0.97$ for the estimate of the part of the association explained by $Z$.
+Here $\text{Cov}(\mathbb{E}[X\mid Z],\mathbb{E}[Y\mid Z])$ is the contribution to the covariance explained by $Z$, and thus possibly due to a confounding effect of $Z$, whereas $\mathbb{E}[\text{Cov}(X,Y\mid Z)]$ is the part that is not explained by $Z$ in the context of a linear framework.
+A question for future research is whether by utilizing the proposed conditional profile association one can extend such a separation of the effect that can be attributed to $Z$ and the remaining association that cannot be explained by $Z$ to scenarios where $X$ and $Y$ are random objects and a linear framework does not apply, due to the lack of vector operations in non-Euclidean spaces.
+Considering the mortality example discussed in Section 6.2, where $X_{i}$ and $Y_{i}$ are female and male age-at-death distributions, while $Z_{i}$ is a Euclidean covariate, a preliminary attempt to shed some light on this is the following three-step procedure: first compute the unconditional profile association $\Delta(X,Y)$, using all available data pairs $\{(X_{i},Y_{i})\}_{i}$; second, replace the conditional expectations $\mathbb{E}[X\mid Z]$, $\mathbb{E}[Y\mid Z]$ in the law of total covariance by Fréchet regressions (Petersen and Müller, 2019) to obtain conditional Fréchet means $\mathbb{E}_{\oplus}[X\mid Z]$ and $\mathbb{E}_{\oplus}[Y\mid Z]$ using data $\{(X_{i},Z_{i})\}_{i}$ and $\{(Y_{i},Z_{i})\}_{i}$ respectively; third, compute $\Delta(\mathbb{E}_{\oplus}[X\mid Z],\mathbb{E}_{\oplus}[Y\mid Z])$ by plugging in the results from the Fréchet regression step.
+Applying these considerations for the human mortality data for all countries combined yields $30\Delta(X,Y)=0.36$ for the estimate of the overall profile association that ignores calendar year, but $30\Delta(\mathbb{E}_{\oplus}[X\mid Z],\mathbb{E}_{\oplus}[Y\mid Z])=0.97$ for the estimate of the part of the association explained by $Z$.
 Here we introduce the multiplier $30$ to normalize the association values so they lie in the interval $[0,1]$.
@@
 This research was supported in part by NSF DMS-2310450. We thank Professor Russell Lyons, Indiana University, for important comments.
-Appendix
-Appendix A Metric spaces of strong negative type
-In metric geometry and probability theory, the concept of negative type generalizes the idea of negative definite matrices from linear algebra and serves as a framework for analyzing how distances in a metric space interact with certain functional inequalities.
-Here we give a brief introduction to the definition and major properties of metric spaces of negative type.
-A detailed discussion of the properties of such spaces and additional examples can be found in Lyons (2013).
-A metric space $(\mathcal{M},d)$ is said to be of negative type if, for any finite collection of points $\omega_{1},\ldots,\omega_{n}\in\mathcal{M}$ and real numbers $\alpha_{1},\ldots,\alpha_{n}$ such that $\sum_{i=1}^{n}\alpha_{i}=0$,
-$\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{i}\alpha_{j}d(\omega_{i},\omega_{j})\leq 0.$
-Examples of metric space of negative types include $\mathcal{L}_{p}$ spaces equipped with the metric $d(\omega_{1},\omega_{2})=\|\omega_{1}-\omega_{2}\|_{p}^{\min\{1,p\}}$ for $0<p\leq 2$, spheres with the geodesic distance and all real and complex hyperbolic spaces (Meckes, 2013).
-Importantly, $(\mathcal{M},d)$ is of negative type if and only if there is a Hilbert space $\mathcal{H}$ and an injective map ϕ::italic-$\phi:\mathcal{M}\mapsto\mathcal{H}$ such that for all $\omega,\omega^{\prime}\in\mathcal{M}$, $d(\omega,\omega^{\prime})=\|\phi(\omega)-\phi(\omega^{\prime})\|^{2}$.
-This equivalent definition was discovered by Schoenberg (1937, 1938), and it is the key to prove that distance profiles uniquely characterize the distribution of random objects.
-It also provides the theoretical foundation for numerous other methods in statistics and machine learning.
-A metric space of strong negative type is a concept that additionally relates to the underlying probability measure in the metric measure space and extends the negative type property beyond the consideration of finite collections of points.
-Specifically, a metric space $(\mathcal{M},d)$ is said to be of strong negative type if the condition
-$D(\mu,\nu):=\iint d(\omega,\omega^{\prime})\,d(\mu-\nu)^{2}(\omega,\omega^{\prime})\leq 0$
-holds for all distinct Borel probability measures $\mu$ and $\nu$ on $\mathcal{M}$ with finite first moments, with equality if and only if $\mu=\nu$.
-The functional $D(\mu,\nu)$ is known as the energy distance between $\mu$ and $\nu$, and it generalizes the notion of a metric on the space of probability measures on the metric space $\mathcal{M}$.
-The strong negative type property ensures that the energy distance is zero if and only if $\mu=\nu$, meaning that the metric structure on the space allows for the identification of probability measures based solely on pairwise distances.